### v1

# Multiple rational normal forms in Lie theory

Abstract
We study the decomposition of a generic element $g\in G$ of a connected reductive complex algebraic group $G$ in the form $g=N(g)B(g)\bar{u}N(g)^{-1}$ where $N:G\dashrightarrow\mathcal{N}_{-}$ and $B:G\dashrightarrow\mathcal{B}_{+}$ are rational maps onto a unipotent subgroup $\mathcal{N}_{-}$ and a Borel subgroup $\mathcal{B}_{+}$ opposite to $\mathcal{N}_{-}$, and $\bar{u}$ is a representative of a Weyl group element $u$. We introduce a class of rational Weyl group elements that give rise to such decompositions, and study their various properties.
00footnotetext: 2010 Mathematics Subject Classification. Primary 20G07; Secondary 20F55, 13F60.00footnotetext: Key words and phrases. Normal form, rational Weyl group element, cluster algebra.
1 Introduction
Let $G$ be a connected complex reductive algebraic group, $(\mathcal{B}_{+},\mathcal{B}_{-})$ a pair of opposite Borel subgroups, $\mathcal{N}_{\pm}\subset\mathcal{B}_{\pm}$ their unipotent radicals, $\mathcal{H}:=\mathcal{B}_{+}\cap\mathcal{B}_{-}$ a Cartan subgroup, $W:=N_{G}(\mathcal{H})/\mathcal{H}$ the corresponding Weyl group. The goal of the paper is to address the following problem:
Problem 1.1.
For a given $G$, describe elements $u$ of its Weyl group $W$ for which there exist rational maps $N:G\dashrightarrow\mathcal{N}_{-}$ and $B:G\dashrightarrow\mathcal{B}_{+}$ such that for a generic $g\in G$,
$g=N(g)B(g)\overline{u}N(g)^{-1}$ (1.1)
where $\overline{u}\in N_{G}(\mathcal{H})$ is a representative of $u$.
Throughout the paper, we call an element $u\in W$ a solution of Problem 1.1 if there exist rational maps $B$ and $N$ that give rise to the decomposition (1.1). In the next example, we illustrate the problem in $\operatorname{GL}_{2}(\mathbb{C})$.
Example 1.2.
For $G:=\operatorname{GL}_{2}(\mathbb{C})$, identify the Weyl group with the group of permutation matrices:
$W=\left\{w_{0}:=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},\ I:=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\right\}.$ (1.2)
Set $D:=\{X\in\operatorname{GL}_{2}(\mathbb{C})\ |\ x_{12}\neq 0\}$. Then every $X\in D$ has a decomposition of the form
$X=\begin{bmatrix}1&0\\ \frac{x_{22}}{x_{12}}&1\end{bmatrix}\begin{bmatrix}x_{12}&x_{11}+x_{22}\\ 0&-\frac{\det X}{x_{12}}\end{bmatrix}\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\begin{bmatrix}1&0\\ -\frac{x_{22}}{x_{12}}&1\end{bmatrix};$ (1.3)
therefore, the element $w_{0}$ is a solution of Problem 1.1. On the other hand, let us set
$\Delta:=\Delta(X):=x_{11}^{2}-2x_{11}x_{12}+x_{22}^{2}+4x_{12}x_{21}.$ (1.4)
Then for the identity permutation, $X\in D$ decomposes as
$X=\begin{bmatrix}1&0\\ \frac{-x_{11}\pm\sqrt{\Delta}}{2x_{12}}&1\end{bmatrix}\begin{bmatrix}\frac{1}{2}(x_{11}+x_{22})\pm\frac{1}{2}\sqrt{\Delta}&x_{12}\\ 0&\frac{1}{2}(x_{11}+x_{22})\mp\frac{1}{2}\sqrt{\Delta}\end{bmatrix}\begin{bmatrix}1&0\\ -\frac{-x_{11}\pm\sqrt{\Delta}}{2x_{12}}&1\end{bmatrix}$ (1.5)
Clearly, the identity permutation is not a solution of Problem 1.1.
Our motivation for studying Problem 1.1 originates in the theory of cluster algebras, as defined by S. Fomin and A. Zelevinsky in [6], and in particular, in its interaction with Poisson geometry. For an overview of the subject and its applications, we refer to [5, 8, 9, 14, 16].
In [10], M. Gekhtman, M. Shapiro, and A. Vainshtein constructed a generalized cluster structure in the ring of invariants [^1] We use the notation $\operatorname{Ad}\mathcal{N}_{-}$ to distinguish the invariant ring from the one where $\mathcal{N}_{-}$ acts by left or right translations. So, $p\in\mathbb{C}[\operatorname{GL}_{n}]^{\operatorname{Ad}\mathcal{N}_{-}}$ if and only if $p(ngn^{-1})=p(g)$ for all $g\in G$, $n\in\mathcal{N}_{-}$. $\mathbb{C}[\operatorname{GL}_{n}]^{\operatorname{Ad}\mathcal{N}-}$, where $\mathcal{N}_{-}$ acts on $\operatorname{GL}_{n}$ by conjugation. The initial extended cluster of $\mathbb{C}[\operatorname{GL}_{n}]^{\operatorname{Ad}\mathcal{N}_{-}}$ is a part of the initial extended clusters in $\mathbb{C}[\operatorname{GL}_{n}]$ and $\mathbb{C}[\operatorname{GL}_{n}\times\operatorname{GL}_{n}]$. In the first case, the obtained cluster structure is compatible [^2] A cluster structure is called compatible with a Poisson bracket if in any given extended cluster, the Poisson bracket is quadratic. with the Poisson bracket induced from the Poisson dual $\operatorname{GL}_{n}^{*}$ of $\operatorname{GL}_{n}$, and in the second case, the Poisson structure on $\operatorname{GL}_{n}\times\operatorname{GL}_{n}$ is that of the Drinfeld double of $\operatorname{GL}_{n}$. As we showed in [11, 15], the same initial extended cluster of $\mathbb{C}[\operatorname{GL}_{n}]^{\operatorname{Ad}\mathcal{N}_{-}}$ is a part of the initial extended clusters in $\mathbb{C}[\operatorname{GL}_{n}]$ and $\mathbb{C}[\operatorname{GL}_{n}\times\operatorname{GL}_{n}]$ compatible with the Poisson brackets from the Belavin-Drinfeld class. We refer to [3, 4] for details on Poisson geometry.
The construction of the initial extended cluster in $\mathbb{C}[\operatorname{GL}_{n}]^{\operatorname{Ad}\mathcal{N}_{-}}$ utilizes a solution of Problem 1.1. Let $C:=s_{1}\cdots s_{n-1}$ be a special Coxeter element in type $A_{n-1}$. By [10, Lemma 8.2], there exist rational maps $B:\operatorname{GL}_{n}(\mathbb{C})\dashrightarrow\mathcal{B}_{+}$ and $N:\operatorname{GL}_{n}(\mathbb{C})\dashrightarrow\mathcal{N}_{-}$ such that for a generic $X\in\operatorname{GL}_{n}(\mathbb{C})$,
$X=N(X)B(X)\overline{C}N(X)^{-1}.$ (1.6)
The initial extended cluster comprises $n-1$ isolated frozen variables, which are the characters $\chi^{\omega_{i}}$ corresponding to the fundamental weights $\omega_{1},\ldots,\omega_{n-1}$, and the rest of the variables is obtained from the initial extended cluster of $\mathbb{C}[\operatorname{GL}_{n-1}]^{\mathcal{N}_{-}}$, where $\mathcal{N}_{-}$ acts by multiplication on the right, via the pullback by the map
$\operatorname{GL}_{n}\ni X\mapsto B(X)^{[2,n]}_{[2,n]}\in\operatorname{GL}_{n-1}.$ (1.7)
It is not known how to extend the above construction to other Lie types. In order to construct a generalized cluster structure in $\mathbb{C}[G]^{\operatorname{Ad}\mathcal{N}_{-}}$ (and in related Poisson varieties), we are thus led to the study of Problem 1.1, which we also find to be interesting in its own right.
Let us briefly explain our results. We call an element $u\in W$ of a Weyl group $W$ rational if a certain graph $\Gamma(u)$ associated with $u$ is acyclic (see Definition 2.3). In [10, Lemma 8.3], it was shown that the longest Weyl group element $w_{0}\in W$ is a solution of Problem 1.1 in any Lie type. We develop a recursive procedure that constructs a pair of maps $(B,N)$ for $u$ from the pair of maps for $w_{0}$ (or, more generally, from any other solution of Problem 1.1). If $u$ is rational, then the recursive procedure ends in a finite number of steps, and thus one obtains the decomposition (1.1) for $u$. We also prove a partial converse: if $u$ is a sufficiently long element and is a solution of Problem 1.1, then $u$ is rational (see Theorem 2.4). We conjecture that the class of solutions of Problem 1.1 is precisely the class of rational Weyl group elements.
We then define a rationality graph $\Gamma(W)$ whose vertices are the rational Weyl group elements of $W$, and two vertices $u,v\in W$ are connected by an edge iff there is a simple reflection $s$ such that $u=sv$. Surprisingly, the graph $\Gamma(W)$ is connected, and it contains more than one vertex if and only if $w_{0}$ induces a nontrivial automorphism of the Dynkin diagram; that is, only in type $A_{r}$, $r\geq 2$, $D_{r}$ for $r$ odd, and $E_{6}$ (see Theorem 2.11; for examples of $\Gamma(W)$, see Appendix A).
In the context of constructing a generalized cluster structure in $\mathbb{C}[\operatorname{SL}_{n}]^{\operatorname{Ad}\mathcal{N}_{-}}$, the special Coxeter element $C$ and its inverse $C^{-1}$ exhibit the following properties: 1) The two elements are the only rational Coxeter elements; 2) The two elements are the only vertices of valency $1$ in $\Gamma(W)$. However, no Coxeter element in any other Lie type is rational; nevertheless, in type $D_{r}$ for $r$ odd, the rationality graph $\Gamma(W)$ also contains two elements of valency $1$, and this similarity leads us to believe that the two elements play a role in the construction of a generalized cluster structure in $\mathbb{C}[\operatorname{Spin}(2r)]^{\operatorname{Ad}\mathcal{N}_{-}}$. We will study this question in our next paper.
We give precise definitions and statements of our main results in the next section. In sections 3 and 4 we provide the proofs.
Acknowledgements.
At various stages of this work, we benefited from discussions with Benjamin Enriquez, Michael Gekhtman, Antoine de Saint Germain, Igor Krylov, Jiang-Hua Lu, Hugh Thomas and Anderson Vera. This research was supported by the Institute for Basic Science (IBS-R003-D1).
2 Main results
2.1 Solutions of Problem 1.1
Let $G$ be a connected complex reductive algebraic group and $\mathfrak{g}$ be its Lie algebra. Denote by $\Pi$ the set of roots of $\mathfrak{g}$, $\Delta$ a set of simple roots, $\Pi_{+}$ and $\Pi_{-}$ the sets of positive and negative roots. We let $\mathcal{B}_{+}$ and $\mathcal{B}_{-}$ to be the corresponding Borel subgroups of $G$, $\mathcal{N}_{\pm}\subseteq\mathcal{B}_{\pm}$ their unipotent radicals, $\mathcal{H}:=\mathcal{B}_{+}\cap\mathcal{B}_{-}$ the corresponding Cartan subgroup, $W:=N_{G}(\mathcal{H})/\mathcal{H}$ the Weyl group. For a subset $A\subseteq\Pi_{+}$, set
$\operatorname{Adj}(A):=\{\alpha\in\Pi_{+}\ |\ \exists\beta\in\Pi_{+}:\ \alpha\leq\beta\}.$ (2.1)
Definition 2.1.
The $\nu$-sequence of an element $u\in W$ is the sequence $\{\nu^{k}(u)\}_{k\geq 0}$ of subsets of $\Pi_{+}$ defined as
$\nu^{0}(u):=u(\Pi_{+})\cap\Pi_{+},\ \ \nu^{k}(u):=u(\operatorname{Adj}\nu^{k-1}(u))\cap\Pi_{+},\ \ k\geq 1.$ (2.2)
In Proposition 3.2, we show that the sequence $\{\nu^{k}(u)\}_{k\geq 0}$ is descending and stabilizes after a finite number of terms. Thus we can make sense of the limit
$\nu(u):=\lim_{k\rightarrow\infty}\nu^{k}(u);$ (2.3)
that is, if $\nu^{m}(u)=\nu^{m+1}(u)=\cdots$ for some $m\geq 1$, then $\nu(u)=\nu^{m}(u)$.
Definition 2.2.
The rationality graph $\Gamma(u)$ of an element $u\in W$ is an oriented graph constructed as follows:
  • The vertex set of $\Gamma(u)$ is given by the set $\nu^{0}(u)$;
  • Two roots $\alpha,\beta\in\nu^{0}(u)$ are connected by an edge $\alpha\rightarrow\beta$ if and only if $u^{-1}(\alpha)\leq\beta$.
Definition 2.3.
An element $u\in W$ is called rational if one of the following equivalent conditions is satisfied:
  1. i)
    The set $\nu(u)$ is empty;
  2. ii)
    The graph $\Gamma(u)$ has no cycles.
We prove the equivalence of the above conditions in Proposition 3.3. In Proposition 3.5 we show that $\Gamma(u)$ contains a loop if and only if $\alpha\leq u(\alpha)$ for some $\alpha\in\Pi_{+}$. As a consequence, a rational element does not fix any root. Another consequence is that if the highest root is a vertex of $\Gamma(u)$, then $u$ is not rational (see Proposition 3.4).
Theorem 2.4.
For any rational element $u\in W$ and its representative $\overline{u}\in N_{G}(\mathcal{H})$, there exist rational maps $B:G\dashrightarrow\mathcal{B}_{+}$ and $N:G\dashrightarrow\mathcal{N}_{-}$ such that for a generic $g\in G$
$g=N(g)B(g)\overline{u}N(g)^{-1}.$ (2.4)
Conversely, if $u\in W$ is any element with $\ell(u)\geq\ell(w_{0})-2$ and the decomposition (2.4) exists for a generic $g\in G$ and some rational maps $B$ and $N$, then $u$ is rational.
In other words, any rational element $u\in W$ is a solution of Problem 1.1, and conversely, any solution $u$ of Problem 1.1 with $\ell(u)\geq\ell(w_{0})-2$ is rational. We conjecture that the solutions of Problem 1.1 are precisely the rational Weyl group elements.
The main part of Theorem 2.4 rests on the following construction, which we precede with a technical definition.
Definition 2.5.
Given $u,v\in W$, the $\nu$-sequence of $u$ relative to $v$ is the sequence $\{\nu^{k}(u\,|\,v)\}_{k\geq 0}$ of subsets of $\Pi_{+}$ defined as follows:
$\nu^{0}(u\,|\,v):=uv^{-1}(\Pi_{-})\cap\Pi_{+};$ (2.5)
$\nu^{k}(u\,|\,v):=u(\operatorname{Adj}\nu^{k-1}(u\,|\,v))\cap\Pi_{+}.$ (2.6)
If the $\nu$-sequence of $u$ relative some $v$ stabilizes after a finitely many terms, we set
$\nu(u\,|\,v):=\lim_{k\rightarrow\infty}\nu^{k}(u\,|\,v).$ (2.7)
Clearly, the $\nu$-sequence of $u$ (see Definition 2.1) is the $\nu$-sequence of $u$ relative to $w_{0}$.
For a generic element $g\in G$, denote by $g_{-}\in\mathcal{N}_{-}$ and $g_{\oplus}\in\mathcal{B}_{+}$ the unique elements such that $g=g_{-}g_{\oplus}$ (the Gaussian decomposition). Let us a fix any solution $v\in W$ of Problem 1.1, and let us denote by $\tilde{B}:G\dashrightarrow\mathcal{B}_{+}$ and $\tilde{N}:G\dashrightarrow\mathcal{N}_{-}$ the rational maps such that for a generic $g\in G$,
$g=\tilde{N}(g)\tilde{B}(g)\overline{v}\tilde{N}(g)^{-1}$ (2.8)
for some representative $\overline{v}$ of $v$. For any other element $u\in W$, construct a sequence of rational maps $\{P_{k}:G\dashrightarrow G\}_{k\geq 0}$ via
$P_{0}(g):=\tilde{B}(g)\overline{v}\,\overline{u}^{-1},$ (2.9)
$P_{k}(g):=[P_{k-1}(g)]_{\oplus}\overline{u}[P_{k-1}(g)]_{-}\overline{u}^{-1},\ \ k\geq 1;$ (2.10)
and a sequence of rational maps $\{N_{k}:G\dashrightarrow\mathcal{N}_{-}\}_{k\geq 0}$ as
$N_{k}(g):=\tilde{N}(g)\tilde{N}(P_{k}(g)\overline{u})^{-1},k\geq 0.$ (2.11)
Proposition 2.6.
For any solution $v\in W$ of Problem 1.1 and any $u\in W$ such that $\nu(u\,|\,v)=\emptyset$, $u$ is also a solution of Problem 1.1. More precisely, for a given representative $\overline{u}$, there is $m\geq 0$ such that for a generic $g\in G$,
$B(g):=P_{m}(g)=P_{m+1}(g)=\ldots\in\mathcal{B}_{+};$ (2.12)
$N(g):=N_{m}(g)=N_{m+1}(g)=\ldots\in\mathcal{N}_{-},$ (2.13)
and such that
$g=N(g)B(g)\overline{u}N(g)^{-1}.$ (2.14)
It was shown in [10, Lemma 8.3] that the longest Weyl group element $w_{0}$ is a solution of Problem 1.1. More precisely, let us set
$N_{0}(g)=[g\overline{w_{0}}^{-1}]_{-};$ (2.15)
$B_{0}(g)=[g\overline{w_{0}}^{-1}]_{\oplus}\overline{w_{0}}([g\overline{w_{0}}^{-1}]_{-})\overline{w_{0}}^{-1}.$ (2.16)
It is straightforward to verify that for a generic $g\in G$,
$g=N_{0}(g)B_{0}(g)\overline{w_{0}}N_{0}(g)^{-1}.$ (2.17)
One direction of Theorem 2.4 is thus obtained as a corollary of Proposition 2.6 by setting $v:=w_{0}$.
To test whether an element $u\in W$ is not a solution of Problem 1.1, we have the following methods: 1) The contrapositive to Proposition 2.6 (see Corollary 3.21; a more practical statement is Corollary 3.22); 2) Determining whether the generic fiber of the restriction $B_{0}|_{\mathcal{B}_{+}\overline{u}}:\mathcal{B}_{+}\overline{u}\rightarrow\mathcal{B}_{+}$ is of cardinality $>1$ (see Lemma 3.24). The statements needed for proving the partial converse to Theorem 2.11 can be summarized as follows.
Proposition 2.7.
Let $u\in W$ be any element and $\alpha\in\Delta$ be a simple root.
  1. 1)
    If $u$ is not a solution of Problem 1.1 and $u(\alpha)\in\Pi_{-}$, then $s_{\alpha}u$ is not a solution of Problem 1.1.
  2. 2)
    If $\Gamma(u)$ has a cycle consisting of pairwise orthogonal simple roots, then $u$ is not a solution of Problem 1.1. In particular, if $u$ fixes a simple root, it is not a solution.
  3. 3)
    Assume there exist $\alpha_{i},\alpha_{j}\in\Delta$ such that $a_{ij}=a_{ji}=-1$ (Cartan integers), with $\alpha_{i},\alpha_{i}+\alpha_{j}\in\nu^{0}(u)$, and $\Gamma(u)$ contains a $2$-cycle on $\alpha_{i}$ and $\alpha_{i}+\alpha_{j}$. Then $u$ is not a solution of Problem 1.1.
The next results concern general properties of the maps $B$ and $N$. For a given $u\in W$ and its representative $\overline{u}$, define a rational map $\mathcal{T}_{\overline{u}}:\mathcal{N}_{-}\times\mathcal{B}_{+}\rightarrow G$ via
$\mathcal{T}_{\overline{u}}(n,b)=nb\overline{u}n^{-1},\ \ n\in\mathcal{N}_{-},\ b\in\mathcal{B}_{+},$ (2.18)
Proposition 2.8.
An element $u\in W$ is a solution of Problem 1.1 if and only if the regular map $\mathcal{T}_{\overline{u}}$ is birational.
Proposition 2.9.
Let $u\in W$ be a solution of Problem 1.1. Then the corresponding rational maps $B:G\dashrightarrow\mathcal{B}_{+}$ and $N:G\dashrightarrow\mathcal{N}_{-}$ satisfy the following properties:
  1. 1)
    Both maps $B$ and $N$ are dominant and unique;
  2. 2)
    For a generic $g\in G$ and $n\in\mathcal{N}_{-}$,
    $B(ngn^{-1})=B(g);$ (2.19)
    $N(ngn^{-1})=nN(g).$ (2.20)
  3. 3)
    The rational map $N$ is determined by $B$ via the equation
    $N(g)=N_{0}(g)N_{0}(B(g)\overline{u})^{-1}=[g\overline{w_{0}}^{-1}]_{-}([B(g)\overline{u}\,\overline{w_{0}}^{-1}]_{-})^{-1}.$ (2.21)
    for any representative $\overline{w_{0}}$ of $w_{0}$.
In particular, in the context of Proposition 2.6, if the sequence $\{N_{k}:G\dashrightarrow\mathcal{N}_{-}\}$ given by (2.11) stabilizes, its limit is given by (2.21).
2.2 Combinatorics of rational Weyl group elements
For this subsection, let us assume that the root system of $G$ is indecomposable.
Definition 2.10.
The rationality graph $\Gamma(W)$ of $W$ is the graph constructed as follows:
  • The vertices of $\Gamma(W)$ are the rational Weyl group elements;
  • Two vertices $u$ and $v$ are connected by an edge if and only if there exists a simple reflection $s_{\alpha}$, $\alpha\in\Delta$, such that $u=s_{\alpha}v$.
If we know that $W$ is of some specific Lie type (say, $A_{r}$), then we write $\Gamma(A_{r})$ instead of $\Gamma(W)$. Some examples of rationality graphs are given in Appendix A.
Theorem 2.11.
The rationality graph $\Gamma(W)$ is connected and carries a $\mathbb{Z}_{2}$-symmetry centered at the vertex $w_{0}$. Moreover, the graph $\Gamma(W)$ has more than one vertex if and only if $w_{0}\neq-\operatorname{id}$.
In other words, there exist rational elements $u\in W$ other than $w_{0}$ if and only if the Lie type of $W$ is either $A_{r}$ for $r\geq 2$, or $D_{r}$ for $r$ odd, or $E_{6}$. Our proof is constructive: for every rational $u\in W$, we find an explicit path to $w_{0}$.
For type $A_{r}$, define the following special Coxeter element:
$C:=s_{1}s_{2}\cdots s_{r}.$ (2.22)
Proposition 2.12.
A Coxeter element $c\in W$ is rational if and only if the Lie type of $W$ is $A_{r}$, $r\geq 1$, and $c\in\{C,C^{-1}\}$. In this case, $c$ has valency $1$ in $\Gamma(A_{r})$.
As we explained in the inroduction, the special Coxeter element (2.22) plays a role in the construction of a generalized cluster structure in the ring of invariants $\mathbb{C}[\operatorname{SL}_{n}]^{\operatorname{Ad}\mathcal{N}_{-}}$ (that is, regular functions invariant under the action $(n,g)\mapsto ngn^{-1}$, $n\in\mathcal{N}_{-}$, $g\in G$). Although Coxeter elements are not rational in type $D_{r}$ for $r$ odd, there is at least the following similarity:
Proposition 2.13.
In type $D_{r}$ for $r\geq 5$ odd, the rationality graph $\Gamma(D_{r})$ contains at least $2$ vertices of valency $1$.
We give an explicit construction of the two elements of valency $1$ in type $D_{r}$ (see Proposition 4.14). In type $E_{6}$, however, there are no rational elements of valency $1$; instead, there are four elements of valency $2$. We expect that both Proposition 2.12 and Proposition 2.13 can be strengthened: the rationality graphs $\Gamma(A_{r})$, $r\geq 2$, and $\Gamma(D_{r})$ for $r$ odd have exactly $2$ vertices of valency $1$.
If our main conjecture holds, the neighborhood of a given vertex $u\in W$ in $\Gamma(W)$ consists of those $s_{\alpha}u$ for which $\nu(s_{\alpha}u\,|\,u)=\emptyset$. In other words, if $u$ is rational, then $s_{\alpha}u$ is rational if and only if $\nu(s_{\alpha}u\,|\,u)=\emptyset$.
In Tables 1 and 2, we present the numbers of rational Weyl group elements in types with nontrivial $w_{0}$. We do not know if there is a closed formula for type $A$; in type $D_{r}$ for $r$ odd, we conjecture that the number of rational Weyl group elements is equal to $12^{r}-$ 1.
Type $A_{1}$ $A_{2}$ $A_{3}$ $A_{4}$ $A_{5}$ $A_{6}$ $A_{7}$ $A_{8}$ $A_{9}$ $A_{10}$
Order of $W$ $2$ $6$ $24$ $120$ $720$ 5 04050405\,0405 040 40 3204032040\,32040 320 362 $80362 8$ 80036288003\,628\,8003 628 800 39 916 8003991680039\,916\,80039 916 800
Rational $u$ $1$ $3$ $7$ $25$ $89$ $379$ 1 67916791\,6791 679 8 28982898\,2898 289 42 5614256142\,56142 561 236 099236099236\,099236 099
Table 1: Count of rational Weyl group elements in type $A_{r}$ up to $r=10$.
Type $D_{5}$ $D_{7}$ $D_{9}$ $D_{11}$ $D_{13}$ $D_{15}$ $E_{6}$
Order of $W$ 1 92019201\,9201 920 322 560322560322\,560322 560 92 897 2809289728092\,897\,28092 897 280 $11!\cdot 2^{10}$ $13!\cdot 2^{12}$ $15!\cdot 2^{14}$ 51 8405184051\,84051 840
Rational $u$ $31$ $127$ $511$ 2 04720472\,0472 047 8 19181918\,1918 191 32 7673276732\,76732 767 $397$
Table 2: Count of rational Weyl group elements in types $D_{r}$, for $r\leq 15$ odd and $E_{6}$.
Remark 2.14.
In the definition of $\Gamma(W)$, we join two rational Weyl group elements $u,v\in W$ iff there is a simple reflection $s$ such that $u=sv$. Alternatively, one could join two elements by an edge iff $u=vs$. However, the resulting graph is not connected unless it contains only $1$ vertex or the Lie type is $A_{1}$ or $A_{2}$.
Remark 2.15.
It is interesting to see whether the notion of a rational Weyl group element can be extended to the non-crystallographic root systems of types $I_{2}(m)$ ($m\notin\{3,4,6\}$), $H_{3}$ and $H_{4}$. For type $I_{2}(m)$, we generate a root system from the generalized Cartan matrix of the form [^3] This choice of the Cartan matrix for $I_{2}(m)$ includes the cases of $C_{2}$ and $G_{2}$ types; however, if one generates a root system of type $I_{2}(m)$ from a symmetric Cartan matrix, the results are not as nice (for the symmetric, for instance, $\Gamma(I_{2}(6))$ has $3$ vertices, which is misaligned with the result for $\Gamma(G_{2})$, which has $1$ vertex; nevertheless, the results for $I_{2}(5)$ and $H_{2}$ agree whether the matrix is symmetric or not).
$\begin{bmatrix}2&-4\cos^{2}(\pi/m)\\ -1&2\end{bmatrix}$ (2.23)
For types $H_{3}$ and $H_{4}$, we choose a symmetric generalized Cartan matrix as in [13, Section 2.13]. The only issue for testing the theory is the choice of a poset. If one defines a poset on the set of roots as
$\alpha\leq\beta\ \ \Leftrightarrow\ \ \beta-\alpha\in\mathbb{R}_{\geq 0}\Delta,$ (2.24)
then the results are as follows. The rationality graph $\Gamma(I_{2}(m))$ is connected, and it contains $1$ vertex if $m$ is even and $3$ vertices if $m$ is odd; the graph $\Gamma(H_{3})$ is not connected and contains $19$ vertices; the graph $\Gamma(H_{4})$ is also not connected and contains 12 4401244012\,44012 440 vertices. Another poset for $I_{2}(m)$ and $H_{3}$ was suggested by D. Armstrong in [1, Section 5.4.1]. We verified that if one induces the Armstrong’s poset in $H_{3}$ from $D_{6}$, then $\Gamma(H_{3})$ contains $4$ vertices; however, the graph is not connected.
3 Construction of solutions
In this section, we provide proofs of the results from Section 2.1.
3.1 Rationality graph $\Gamma(u)$ and $\nu$-sequence
In this subsection, for a given $u\in W$, we study properties of the $\nu$-sequence of $u$ (see Definition 2.1), properties of its rationality graph $\Gamma(u)$ (see Definition 2.2), and the relation between $\Gamma(u)$ and the $\nu$-sequence. We also prove the equivalence of conditions i and ii of Definition 2.3 (see Proposition 3.3).
Lemma 3.1.
For any $k\geq 0$, a root $\alpha\in\nu^{0}(u)$ belongs to $\nu^{k}(u)$ if and only if there is a path in $\Gamma(u)$ of length $k$ that starts at $\alpha$.
Note that we allow loops in $\Gamma(u)$, so a path can traverse the same vertex multiple times.
Proof.
Let us prove the statement by induction. For $k=0$, the vertex set of $\Gamma(u)$ is the same as the set of paths of length 00, and it is equal to $\nu^{0}(u)$. Assume that the statement holds for $k-1$, $k\geq 1$. By equation (2.2), a root $\alpha\in\nu^{0}(u)$ belongs to $\nu^{k}(u)$ if and only if there is a root $\beta\in\nu^{k-1}(u)$ such that $u^{-1}(\alpha)\leq\beta$; that is, there is an arrow $\alpha\rightarrow\beta$ in the graph $\Gamma(u)$. But then, there is a path of length $k-1$ that starts at $\beta$; thus, there is a path of length $k$ that starts at $\alpha$. ∎
Proposition 3.2.
The sequence $\{\nu^{k}(u)\}_{k\geq 0}$ is descending and stabilizes after a finite number of terms.
Proof.
By Lemma 3.1, a root $\alpha\in\nu^{0}(u)$ belongs to $\nu^{k}(u)$ if and only if there is a path of length $k$ that starts at $\alpha$; hence, if $k\geq 1$ and $\alpha\in\nu^{k}(u)$, there is also a path of length $k-1$ that starts at $\alpha$, and thus $\alpha\in\nu^{k-1}(u)$. Since the elements of the sequence are subsets of a finite set, and since the sequence is descending, it stabilizes after a finite number of terms. ∎
The result of Proposition 3.2 allows us to define
$\nu(u):=\lim_{k\rightarrow\infty}\nu^{k}(u).$ (3.1)
As a consequence, the set $\nu(u)$ satisfies the following equation:
$\nu(u)=u(\operatorname{Adj}\nu(u))\cap\Pi_{+}.$ (3.2)
The following result shows that the two definitions of rationality of an element $u\in W$ are equivalent (see Definition 2.3).
Proposition 3.3.
A positive root $\alpha$ belongs to $\nu(u)$ if and only if there is a cycle in $\Gamma(u)$ that passes through $\alpha$. In particular, the graph $\Gamma(u)$ contains a cycle if and only if $\nu(u)$ is nonempty.
Proof.
If $\nu(u)$ is nonempty, we can find paths in $\Gamma(u)$ of arbitrary length, and since $\Gamma(u)$ has a finite number of vertices and is oriented, it contains a cycle. Conversely, if $\Gamma(u)$ contains a cycle, by Lemma 3.1, for any $k\geq 1$, the set $\nu^{k}(u)$ is nonempty, and thus $\nu(u)$ is nonempty as well. ∎
Proposition 3.4.
Let $\theta$ be the highest root. If $\theta\in\nu^{0}(u)$, then $u$ is not rational.
Proof.
Indeed, if $\theta\in\nu^{0}(u)$, then there is a loop in $\Gamma(u)$ at $\theta$, for $u^{-1}(\theta)\leq\theta$. ∎
Proposition 3.5.
The graph $\Gamma(u)$ contains a loop if and only if there is a root $\alpha\in\Pi_{+}$ such that $\alpha\leq u(\alpha)$. In particular, if $u$ has a fixed point, then $\Gamma(u)$ contains a loop.
Proof.
Indeed, $\Gamma(u)$ has a loop if and only if there is $\beta\in\nu^{0}(u)$ such that $u^{-1}(\beta)\leq\beta$. For any $\beta\in\nu^{0}(u)$, there is $\alpha\in\Pi_{+}$ such that $\beta=u(\alpha)$; thus, $u^{-1}(\beta)\leq\beta$ if and only if $\alpha\leq u(\alpha)$. ∎
For the following corollary, identify the Weyl group of type $A_{n-1}$ with $S_{n}$. Although we are not going to use the next result, we find it curious enough to record here.
Corollary 3.6.
For an element $u\in S_{n}$, the following statements hold:
  1. 1)
    If for some $k\in\{1,\ldots,n-1\}$, $u$ has the cycle (1 $(1\,2\,\ldots\,k)$ in its cycle decomposition, then $u$ is not rational;
  2. 2)
    If for some $k\in\{2,\ldots,n\}$, $u$ has the cycle $(k\,k+1\,\ldots\,n)$ in its cycle decomposition, then $u$ is not rational.
Proof.
  1. 1)
    Assume $u$ has the cycle (1 $(1\,2\,\ldots\,k)$ in its cycle decomposition, and let $j:=u(k+1)$. Then $u(\alpha_{k})=\alpha_{1}+\cdots+\alpha_{j-1}\geq\alpha_{k}$. By Proposition 3.5, the graph $\Gamma(u)$ contains a loop, and thus $u$ is not rational.
  2. 2)
    Similarly, if $u$ has the cycle $(k\,k+1\,\ldots\,n)$ in its cycle decomposition, then $u(\alpha_{k-1})\geq\alpha_{k-1}$, and therefore, $u$ is not rational. ∎
Remark 3.7.
[…]
Remark 3.8.
The set $\nu^{0}(u)$ can be computed as follows. Write $u=vw_{0}$ for some $v\in W$, and decompose $v$ into its reduced decomposition $v=s_{i_{1}}s_{i_{2}}\cdots s_{i_{m}}$. Set
$\beta_{1}:=\alpha_{i_{1}},\ \ \beta_{k}:=s_{i_{1}}s_{i_{2}}\cdots s_{i_{k-1}}(\alpha_{k}),\ \ k\geq 2.$ (3.3)
Then the set $\nu^{0}(u)$ is given by
$\nu^{0}(u)=\{\beta_{1},\beta_{2},\ldots,\beta_{m}\}.$ (3.4)
Indeed, $\nu^{0}(u)$ is the set of positive roots $\alpha$ such that $u^{-1}(\alpha)$ is positive; alternatively, it is the set of positive roots $\alpha$ such that $v^{-1}(\alpha)$ is negative, and the computation of the latter set via reduced decompositions is well-known. More generally, for any $w\in W$, the set $\nu^{0}(u\,|\,w)$ can be computed in the same way by considering a reduced decomposition of $uw^{-1}$.
The last results of the subsection concern relative $\nu$-sequences (see Definition 2.5).
Proposition 3.9.
For any $u,v\in W$, set $w:=uv^{-1}$, and decompose $w$ into its reduced decomposition as $w=s_{i_{1}}\cdots s_{i_{m}}$. Then
$\{\alpha_{i_{1}},\ldots,\alpha_{i_{m}}\}\subseteq\operatorname{Adj}\nu^{0}(u\,|\,v)\subseteq\Pi_{+}\cap\mathbb{Z}_{+}\{\alpha_{i_{1}},\ldots,\alpha_{i_{m}}\}.$ (3.5)
Proof.
Write $\nu^{0}(u\,|\,v)=\{\beta_{1},\ldots,\beta_{m}\}$ where $\beta_{i}$’s are given by (3.3). We proceed by induction on $\ell(w)$. If $\ell(w)=1$, then $\operatorname{Adj}\nu^{0}(u\,|\,v)=\{\alpha_{i_{1}}\}$. Assuming that $\ell(w)=m-1$, observe that there are two options: if $i_{m}=i_{k}$ for some $k\leq m-1$, then $\alpha_{i_{m}}\in\operatorname{Adj}\nu^{0}(u\,|\,v)$ by the inductive assumption; otherwise, $\beta_{m}=s_{i_{1}}\cdots s_{i_{m-1}}(\alpha_{i_{m}})$, and since $i_{k}\neq i_{m}$ for all $k\leq m-1$, we see that $\alpha_{i_{m}}\leq\beta_{m}$. ∎
Proposition 3.10.
If $u\in W$ is rational, then for any $v\in W$, $\nu(u\,|\,v)=\emptyset$.
Proof.
Since $\nu^{0}(u\,|\,v)\subseteq\Pi_{+}$, we see that $\nu^{1}(u\,|\,v)\subseteq\nu^{0}(u)$. Reasoning inductively, we conclude that $\nu^{k+1}(u\,|\,v)\subseteq\nu^{k}(u)$; thus, if $u$ is rational, $\nu(u\,|\,v)=\emptyset$ for any $v\in W$. ∎
Remark 3.11.
For a given $u\in W$, the $\nu$-sequence of $u$ is descending (see Proposition 3.2); however, relative $\nu$-sequences may not be descending. Consider, for example, type $A_{3}$ with $u:=s_{3}s_{1}s_{2}$ and $v:=s_{1}s_{2}$. One finds that $\nu^{0}(u\,|\,v)=\{\alpha_{3}\}$ and
$\nu^{k}(u\,|\,v)=\begin{cases}\{\alpha_{1}+\alpha_{2}\}&\text{if}\ k\ \text{isodd};\\ \{\alpha_{2}+\alpha_{3}\}&\text{if}\ k\ \text{is even.}\end{cases}$ (3.6)
Remark 3.12.
Rationality of an element $u\in W$ can be linked to the nilpotency of the operator $T_{b,\overline{u}}$ described as follows. Let $\mathfrak{n}_{\pm}$ be the Lie algebra of $\mathcal{N}_{\pm}$, $\pi_{<}:\mathfrak{g}\rightarrow\mathfrak{n}_{-}$ be the orthogonal projection. For an element $b\in\mathcal{B}_{+}$, a given $u\in W$ and its representative $\overline{u}\in N_{G}(\mathcal{H})$, define the operator $T_{b,\overline{u}}:\mathfrak{n}_{+}\rightarrow\mathfrak{n}_{+}$ as
$T_{b,\overline{u}}(x):=\pi_{<}\operatorname{Ad}_{b\overline{u}}(x),\ \ x\in\mathfrak{n}_{-}.$ (3.7)
It is not difficult to see that the image of $(T_{b,\overline{u}})^{i}$ satisfies the equation
$(T_{b,\overline{u}})^{i}(\mathfrak{n}_{-})\subseteq\bigoplus_{\beta\in\operatorname{Adj}\nu^{i-1}(u)}\mathfrak{g}_{-\beta},\ \ i\geq 1;$ (3.8)
therefore, if $u$ is rational, $T_{b,\overline{u}}$ is nilpotent. We expect that the converse is also true: if $T_{b,\overline{u}}$ is nilpotent for a generic $b\in\mathcal{B}_{+}$, then $u$ is rational. However, we find no use of this map, except that it appears in the formula for the differential of the map $B$.
3.2 Proofs of Proposition 2.8 and 2.9 (General properties)
We formulated Problem 1.1 as the problem of finding a pair of rational maps $B:G\dashrightarrow\mathcal{B}_{+}$ and $N:G\dashrightarrow\mathcal{N}_{-}$. In this subsection, we show that, equivalently, we could formulate Problem 1.1 as the problem of birationality of the map
$\mathcal{T}_{\overline{u}}:\mathcal{N}_{-}\times\mathcal{B}_{+}\rightarrow G,\ \ \mathcal{T}_{\overline{u}}(n,b)=nb\overline{u}n^{-1},\ \ n\in\mathcal{N}_{-},\ b\in\mathcal{B}_{+},$ (3.9)
for a fixed representative $\overline{u}$ of $u$. This equivalence follows from a simple observation in algebraic geometry (see Lemma 3.13). As a consequence, we derive Proposition 2.9, which concerns some general properties of the maps $B$ and $N$. The statements of Proposition 2.9 are split into three statements in this subsection: Corollary 3.14, Corollary 3.15 and Proposition 3.16.
Lemma 3.13.
Let $V$ and $W$ be complex projective varieties of the same dimension $n:=\dim V=\dim W$, and let $f:V\dashrightarrow W$ be a rational map. Assume there exists a rational map $s:W\dashrightarrow V$ such that $fs=\operatorname{id}$. Then $f$ is birational.
Proof.
The condition $fs=\operatorname{id}$ implies that $s$ is generically one-to-one and has an injective differential. Together with the assumption that both $V$ and $W$ are projective, by [12, Corollary 14.10], we conclude that $s$ is birational. Thus $f$ is birational as well. ∎
Proof of Proposition 2.8.
Note that both the domain and the codomain of $\mathcal{T}_{\overline{u}}$ are rational varieties. The product of the maps $N$ and $B$ yields a rational section of $\mathcal{T}_{\overline{u}}$, and therefore, by Lemma 3.13, the map $\mathcal{T}_{\overline{u}}$ is birational. ∎
Corollary 3.14.
For any solution $u\in W$ of Problem 1.1, the rational maps $B:G\dashrightarrow\mathcal{B}_{+}$ and $N:G\dashrightarrow\mathcal{N}_{-}$ are dominant.
Proof.
Indeed, since the product map $N\times B$ is inverse to $\mathcal{T}_{\overline{u}}$, its components $N$ and $B$ are dominant maps. ∎
Corollary 3.15.
For any solution $u\in W$ of Problem 1.1, the corresponding rational maps $B$ and $N$ satisfy the following invariance properties: for any $n\in\mathcal{N}_{-}$ and a generic $g\in G$,
$B(ngn^{-1})=B(g);$ (3.10)
$N(ngn^{-1})=nN(g).$ (3.11)
Proof.
Given any $n\in\mathcal{N}_{-}$, define the rational maps $\bar{N}(g):=n^{-1}N(ngn^{-1})$ and $\bar{B}(g):=B(ngn^{-1})$. We see that
$ngn^{-1}=N(ngn^{-1})B(ngn^{-1}){\overline{u}}N(ngn^{-1})^{-1};$ (3.12)
conjugating both sides by $n^{-1}(\cdot)n$, we see that the above equation can be rewritten as
$g=\bar{N}(g)\bar{B}(g){\overline{u}}\bar{N}(g)^{-1}.$ (3.13)
It follows that $\bar{N}\times\bar{B}$ is a rational section of the map $\mathcal{T}_{\overline{u}}$, and since $\mathcal{T}_{\overline{u}}$ is birational, $N(g)=\bar{N}(g)$ and $B(g)=\bar{B}(g)$. ∎
Proposition 3.16.
For any solution $u$ of Problem 1.1 and its any representative ${\overline{u}}$, the rational map $N:G\dashrightarrow\mathcal{N}_{-}$ is given by
$N(g)=N_{0}(g)N_{0}(B(g){\overline{u}})^{-1}=[g\overline{w_{0}}^{-1}]_{-}([B(g)\overline{u}\,\overline{w_{0}}^{-1}]_{-})^{-1},$ (3.14)
where $\overline{w_{0}}$ is any representative of $w_{0}$.
Proof.
Indeed, decompose a generic $g\in G$ relative ${\overline{u}}$, and decompose $B(g){\overline{u}}$ relative $\overline{w_{0}}$ to obtain the equations
$g=N(g)B(g){\overline{u}}N(g)^{-1};$ (3.15)
$B(g){\overline{u}}=N_{0}(B(g){\overline{u}})B_{0}(g)\overline{w_{0}}N_{0}(B(g){\overline{u}})^{-1}.$ (3.16)
Substituting $B(g){\overline{u}}$ from the second equation into the first equation, we see that (3.14) holds. ∎
3.3 Preliminaries on Lie theory
For the next subsections, we set up notation and review some background on Lie theory.
Gaussian decomposition.
For a generic $g\in G$, we denote by $g_{-}\in\mathcal{N}_{-}$ and $g_{\oplus}$ the unique elements such that $g=g_{-}g_{\oplus}$. For any $a\in\mathcal{H}$, $n_{-}\in\mathcal{N}_{-}$ and $n_{+}\in\mathcal{N}_{+}$, the Gaussian decomposition has the following properties:
$(n_{-}g)_{-}=n_{-}g_{-},$ $(gan_{+})_{-}=g_{-},$ $(ag)_{-}=ag_{-}a^{-1};$ (3.17)
$(n_{-}g)_{\oplus}=g_{\oplus},$ $(gan_{+})_{\oplus}=g_{\oplus}an_{+},$ $(ag)_{\oplus}=ag_{\oplus}.$ (3.18)
We use these properties routinely in Section 3.4.
Cartan integers.
We use the following convention on Cartan integers $a_{ij}$:
$a_{ij}:=\langle\alpha_{j},\alpha_{i}^{\vee}\rangle=2\frac{\langle\alpha_{j},\alpha_{i}\rangle}{\langle\alpha_{i},\alpha_{i}\rangle}.$ (3.19)
In this convention, the simple reflections act on simple roots as
$s_{i}(\alpha_{j})=\alpha_{j}-a_{ij}\alpha_{i}.$ (3.20)
Chevalley generators.
Let $\{(e_{i},h_{i},f_{i})\}_{i=1}^{r}$ be a collection of Chevalley generators of $\mathfrak{g}$, where $r$ is the rank of $\mathfrak{g}$ and $e_{i}:=e_{\alpha_{i}}$, $\alpha_{i}\in\Delta$. We complement the Chevalley generators to a basis of $\mathfrak{g}$ by choosing arbitrary nonzero root vectors $e_{\beta}\in\mathfrak{g}_{\beta}$, $\beta\in\Pi$. Define the following $1$-parameter subgroups of $\mathfrak{g}$:
$x_{\beta}:=\exp(te_{\beta}),\ \ t\in\mathbb{C};$ (3.21)
$x_{i}:=x_{\alpha_{i}},\ \ y_{i}:=x_{-\alpha_{i}},\ \ i\in\{1,\ldots,r\}.$ (3.22)
For each $i\in\{1,\ldots,r\}$, define a group homomorphism $\phi_{i}:\operatorname{SL}_{2}(\mathbb{C})\rightarrow G$ via
$\phi_{i}\begin{bmatrix}1&t\\ 0&1\end{bmatrix}:=x_{i}(t),\ \ \phi_{i}\begin{bmatrix}1&0\\ t&1\end{bmatrix}:=y_{i}(t).$ (3.23)
Define the element $t^{h_{i}}\in\mathcal{H}$ via
$t^{h_{i}}:=\phi_{i}\begin{bmatrix}t&0\\ 0&t^{-1}\end{bmatrix}$ (3.24)
The group versions $x_{i}$ and $y_{i}$ of Chevalley generators satisfy various relations that are well-documented in [2, Section 3] and [7, Section 2]; we will need some of them. For any $t_{1},t_{2}\in\mathbb{C}$ and $i\neq j$,
$x_{i}(t_{1})y_{j}(t_{2})=y_{j}(t_{2})x_{i}(t_{1});$ (3.25)
$x_{i}(t_{1})y_{i}(t_{2})=y_{i}\left(\frac{t_{2}}{1+t_{1}t_{2}}\right)(1+t_{1}t_{2})^{h_{i}}x_{i}\left(\frac{t_{1}}{1+t_{1}t_{2}}\right)\ \ \ (t_{1}t_{2}\neq-1);$ (3.26)
if $a_{ij}=0$, then
$x_{i}(t_{1})x_{j}(t_{2})=x_{j}(t_{2})x_{i}(t_{1});$ (3.27)
$y_{i}(t_{1})y_{j}(t_{2})=y_{j}(t_{2})y_{i}(t_{1});$ (3.28)
if $a_{ij}=a_{ji}=-1$ and $t_{1}\neq-t_{3}$, then
$x_{i}(t_{1})x_{j}(t_{2})x_{i}(t_{3})=x_{j}\left(\frac{t_{2}t_{3}}{t_{1}+t_{3}}\right)x_{i}(t_{1}+t_{3})x_{j}\left(\frac{t_{1}t_{2}}{t_{1}+t_{3}}\right);$ (3.29)
$y_{i}(t_{1})y_{j}(t_{2})y_{i}(t_{3})=y_{j}\left(\frac{t_{2}t_{3}}{t_{1}+t_{3}}\right)y_{i}(t_{1}+t_{3})y_{j}\left(\frac{t_{1}t_{2}}{t_{1}+t_{3}}\right).$ (3.30)
For any $\alpha_{i}\in\Delta$ and an element $a:=\exp(h)\in\mathcal{H}$, $h\in\mathfrak{h}$, we define the character $a\mapsto a^{\alpha_{i}}$ via
$a^{\alpha_{i}}:=e^{\alpha_{i}(h)};$ (3.31)
then the generators $x_{i}$ and $y_{i}$ satisfy the following relations:
$ax_{i}(t)=x_{i}(a^{\alpha_{i}}t)a;$ (3.32)
$ay_{i}(t)=y_{i}(a^{-\alpha_{i}}t)a.$ (3.33)
Special representatives.
For a simple reflection $s_{i}$, define $\dot{s}_{i}\in N_{G}(\mathcal{H})$ to be a representative given by
$\dot{s}_{i}:=x_{i}(-1)y_{i}(1)x_{i}(-1).$ (3.34)
It is well-known that the representatives $\dot{s}_{i}$ satisfy braid relations; therefore, $\dot{u}\in N_{G}(\mathcal{H})$ is well-defined. Our statements do not depend on the particular choice of $\overline{u}\in N_{G}(\mathcal{H})$, so we often choose $\dot{u}$.
3.4 Proof of Proposition 2.6 (Parabolic approximations)
Let us recall the setup of Proposition 2.6. Fix any solution $v\in W$ of Problem 1.1 and its representative $\overline{v}$, and denote by $\tilde{B}:G\dashrightarrow\mathcal{B}_{+}$ and $\tilde{N}:G\dashrightarrow\mathcal{N}_{-}$ the rational maps such that for a generic $g\in G$,
$g=\tilde{N}(g)\tilde{B}(g)\overline{v}\tilde{N}(g)^{-1}.$ (3.35)
For any other $u\in W$ and its representative $\overline{u}$, construct a sequence $\{P_{k}:G\dashrightarrow G\}_{k\geq 0}$ of rational maps via
$P_{0}(g):=\tilde{B}(g)\overline{v}\,\overline{u}^{-1};$ (3.36)
$P_{k}(g):=[P_{k-1}(g)]_{\oplus}\overline{u}[P_{k-1}(g)]_{-}\overline{u}^{-1},\ \ k\geq 1.$ (3.37)
and construct a sequence of rational maps $\{N_{k}:G\dashrightarrow\mathcal{N}_{-}\}$ via
$N_{k}(g):=\tilde{N}(g)\tilde{N}(P_{k}(g)\overline{u})^{-1},\ \ k\geq 0.$ (3.38)
Below we show that if the sequence $\{P_{k}:G\dashrightarrow G\}_{k\geq 0}$ stabilizes after a finitely many terms, then $u$ is a solution of Problem 1.1.
Lemma 3.17.
For any $k\geq 1$, the following equation holds:
$[P_{k-1}(g)]_{-}N_{k}(P_{k}(g)\overline{u})=\tilde{N}(P_{k-1}(g)\overline{u}).$ (3.39)
Proof.
Let us recall from Corollary 3.15 that for any $n\in\mathcal{N}_{-}$ and a generic $g\in G$, $\tilde{N}(ngn^{-1})=n\tilde{N}(g)$. From equation (3.37), we see that
$\begin{split}N_{k}(P_{k}(g)\overline{u})=&N_{k}([P_{k-1}(g)]_{\oplus}\overline{u}[P_{k-1}(g)]_{-})=\\ =&\tilde{N}\left(([P_{k-1}(g)]_{-})^{-1}[P_{k-1}(g)]_{-}[P_{k-1}(g)]_{\oplus}\overline{u}[P_{k-1}(g)]_{-}\right)=\\ =&([P_{k-1}(g)]_{-})^{-1}\tilde{N}(P_{k-1}(g)\overline{u}).\end{split}$ (3.40)
Thus equation (3.39) holds. ∎
Proposition 3.18.
For any $k\geq 0$ and a generic $g\in G$, the following equation holds:
$g=N_{k}(g)P_{k}(g)\overline{u}N_{k}(g)^{-1}.$ (3.41)
Proof.
Let us proceed by induction. For $k=0$, the decomposition (3.41) is precisely the decomposition (3.35) for $v$. Assuming that the decomposition exists for $k-1$, we see that
$\begin{split}g=N_{k-1}(g)&P_{k-1}(g)\overline{u}N_{k-1}(g)^{-1}=\\ =&\tilde{N}(g)\tilde{N}(P_{k-1}(g)\overline{u})^{-1}P_{k-1}(g)\overline{u}\tilde{N}(P_{k-1}(g)\overline{u})\tilde{N}(g)^{-1};\end{split}$ (3.42)
using equation (3.39), we can rewrite the latter as
$\begin{split}g=\overbrace{\tilde{N}(g)\tilde{N}(P_{k}(g)\overline{u})^{-1}}^{N_{k}(g)}\underbrace{([P_{k-1}(g)]_{-})^{-1}P_{k-1}(g)}_{[P_{k-1}(g)]_{\oplus}}\overline{u}[P_{k-1}(g)]_{-}\overbrace{\tilde{N}(P_{k}(g)\overline{u})\tilde{N}(g)^{-1}}^{N_{k}(g)^{-1}}=\\ =N_{k}(g)P_{k}(g)\overline{u}N_{k}(g)^{-1}.\end{split}$ (3.43)
Thus equation (3.41) holds. ∎
Corollary 3.19.
Assume that there is $m\geq 0$ such that $P_{m}(g)\in\mathcal{B}_{+}$ for a generic $g$. Then the sequence $\{P_{k}:G\dashrightarrow G\}_{k\geq 0}$ stabilizes; that is,
$P_{m}(g)=P_{m+1}(g)=\cdots$ (3.44)
Moreover, setting $B(g):=P_{m}(g)$ and $N(g):=N_{m}(g)$, for a generic $g\in G$, the following decomposition holds:
$g=N(g)B(g)\overline{u}N(g)^{-1}.$ (3.45)
Proof.
Indeed, if $P_{m}(g)\in\mathcal{B}_{+}$ for a generic $g\in G$, then it follows from (2.10) that $P_{m+1}(g)=P_{m}(g)$, and so on for the next terms. Thus we see that (3.45) is the decomposition (3.41) for $k:=m$. ∎
We now turn to the proof of Proposition 2.6, which states that if $\nu(u\,|\,v)=\emptyset$, then the sequences $\{P_{k}:G\dashrightarrow G\}_{k\geq 0}$ and $\{N_{k}:G\dashrightarrow\mathcal{N}_{-}\}_{k\geq 0}$ stabilize, and hence $u$ is a solution of Problem 1.1.
Proof of Proposition 2.6.
We employ the notation that was set up in Section 3.3. The statement does not depend on the given representatives of $u$ and $v$, so we choose the ones determined by (3.34). By Corollary 3.19, it suffices to show that there is $m\geq 0$ such that for a generic $g\in G$, $P_{m}(g)\in\mathcal{B}_{+}$. Let us set
$\Delta^{k}:=\Delta\cap\operatorname{Adj}\nu^{k}(u\,|\,v).$ (3.46)
Decomposing $uv^{-1}$ into its reduced decomposition $uv^{-1}=s_{i_{1}}\cdots s_{i_{m}}$, we see that, by Proposition 3.9,
$\Delta^{0}=\{\alpha_{i_{1}},\ldots,\alpha_{i_{m}}\}\subseteq\operatorname{Adj}\nu^{0}(u\,|\,v).$ (3.47)
Let $\mathcal{P}_{k}$ be the parabolic subgroup that is generated by $x_{\beta}(t)$, $\beta\in\Pi_{+}\cup(\Pi_{-}\cap\mathbb{Z}\Delta^{k})$. We see that $P_{0}(g)$ is the product of elements from $\mathcal{B}_{+}$ and $y_{i_{j}}(1)$, $j\in\{1,\ldots,m\}$ (here we use formula (3.34) for the specific choices of $\overline{u}$ and $\overline{v}$); hence, $P_{0}(g)\overline{u}\in\mathcal{P}_{0}$ for a generic $g\in G$. Assuming that $P_{k-1}(g)\in\mathcal{P}_{k-1}$, we see that $[P_{k-1}(g)]_{-}$ can be written as the product of some $x_{-\alpha}(t_{\alpha})$ for $\alpha\in\Delta^{k-1}$; therefore, $\overline{u}[P_{k-1}(g)]_{-}\overline{u}^{-1}$ is the product of $x_{-u(\alpha)}(c_{\alpha}t_{\alpha})$ for $\alpha\in\Delta^{k-1}$ and some $c_{\alpha}\in\mathbb{C}$, and if $u(\alpha)>0$, we see that $u(\alpha)\in\nu^{k}(u\,|\,v)$. The latter implies that if $u(\alpha)>0$, then $u(\alpha)\in\mathbb{Z}_{+}\Delta^{k}$, and in particular, $x_{-u(\alpha)}(t)$ can be written as the product of some $x_{-\beta}(t_{\beta})$ for $\beta\in\Delta^{k}$. We conclude that $P_{k}(g)\in\mathcal{P}_{k}$, and since $\nu(u\,|\,v)=\emptyset$, by Corollary 3.19, there is $m\geq 0$ such that $P_{m}(g)\in\mathcal{B}_{+}$. Thus $u$ is a solution of Problem 1.1. ∎
Corollary 3.20.
Any rational Weyl group element is a solution of Problem 1.1.
Proof.
Indeed, for rational Weyl group elements $u\in W$, $\nu(u)=\nu(u\,|\,w_{0})=\emptyset$. ∎
Corollary 3.21.
Assume that $u$ is not a solution of Problem 1.1. Then for any $v\in W$ such that $\nu(u\,|\,v)=\emptyset$, $v$ is not a solution of Problem 1.1 as well.
The next corollary corresponds to Statement 1 of Proposition 2.7.
Corollary 3.22.
Let $u\in W$ be an element such that $u(\alpha)\in\Pi_{-}$ and $u$ is not a solution of Problem 1.1. Then $s_{\alpha}u$ is not a solution of Problem 1.1 as well.
Proof.
Computing the relative $\nu$-sequence, we see that
$\nu^{0}(u\,|\,s_{\alpha}u)=\{\alpha\};$ (3.48)
$\nu^{1}(u\,|\,s_{\alpha}u)=\{u(\alpha)\}\cap\Pi_{+}=\emptyset.$ (3.49)
By Corollary 3.21, $s_{\alpha}u$ is not a solution. ∎
Remark 3.23.
For any $u\in W$, one can also show the invariance of each term of the sequences $\{N_{k}:G\dashrightarrow\mathcal{N}_{-}\}_{k\geq 0}$ and $\{P_{k}:G\dashrightarrow G\}_{k\geq 0}$:
$P_{k}(ngn^{-1})=P_{k}(g);$ (3.50)
$N_{k}(ngn^{-1})=nN_{k}(g).$ (3.51)
3.5 Proof of Proposition 2.7 (Nonexistence of solutions)
In this subsection, we develop methods for showing that an element $u\in W$ is not a solution of Problem 1.1. One such method can be derived from the contrapositive of Proposition 2.6 (see Corollary 3.21 and Corollary 3.22); the other method consists of analyzing the fibers of the pair of maps $(B_{0},N_{0})$ for the solution $w_{0}$. In particular, we prove statements 2 and 3 of Proposition 2.7, which are split into three other statements: Proposition 3.25, Proposition 3.26 and Proposition 3.27. Throughout the subsection, we employ the notation that was set up in Section 3.3.
Lemma 3.24.
Let $u\in W$ be any Weyl group element and $\overline{u}$ be its representative. Assume that for a generic $b\in\mathcal{B}_{+}$, there exists $n:=n(b)\in\mathcal{N}_{-}$ such that $n\neq 1$ and $n^{-1}b\overline{u}n\overline{u}^{-1}\in\mathcal{B}_{+}$. Then $u$ is not a solution of Problem 1.1.
Proof.
Indeed, pick any representative $\overline{w_{0}}$ of $w_{0}$ and the pair of maps $(B_{0},N_{0})$ as in (2.15)-(2.16). Consider the rational map $L_{0}:\mathcal{B}_{+}\overline{u}\dashrightarrow\mathcal{B}_{+}\overline{w_{0}}$ given by $L_{0}(b\overline{u})=B_{0}(b\overline{u})\overline{w_{0}}$. Observe that if $L_{0}(b\overline{u})=L_{0}(b^{\prime}\overline{u})$, then
$b\overline{u}=N_{0}(b\overline{u})L_{0}(b\overline{u})N_{0}(b\overline{u})^{-1},$ (3.52)
$b^{\prime}\overline{u}=N_{0}(b^{\prime}\overline{u})L_{0}(b^{\prime}\overline{u})N_{0}(b^{\prime}\overline{u})^{-1},$ (3.53)
and therefore, by setting $n:=N_{0}(b\overline{u})N_{0}(b^{\prime}\overline{u})^{-1}$, we arrive at the relation
$b^{\prime}=n^{-1}b\overline{u}n\overline{u}^{-1}\in\mathcal{B}_{+}.$ (3.54)
Clearly, $b^{\prime}\neq b$ if and only if $n\neq 1$.
Now, let us show that if $u$ is a solution of Problem 1.1, then the generic fiber of $L_{0}$ is necessarily of size $1$. Let $N:G\dashrightarrow\mathcal{N}_{-}$ and $B:G\dashrightarrow\mathcal{B}_{+}$ be the pair of rational maps that give rise to the rational decomposition with $\overline{u}$. Define the map $L:\mathcal{B}_{+}\overline{w_{0}}\dashrightarrow\mathcal{B}_{+}\overline{u}$ via $L(b\overline{w_{0}})=B(b\overline{w_{0}})\overline{u}$. Then the rational maps $L$ and $L_{0}$ are inverse to each other. Indeed,
$L_{0}(L(b\overline{w_{0}}))=B_{0}(B(b\overline{w_{0}})\overline{u})\overline{w_{0}}=B_{0}(N(b\overline{w_{0}})B(b\overline{w_{0}})\overline{u}N(b\overline{w_{0}})^{-1})\overline{w_{0}}=b\overline{w_{0}}$ (3.55)
where we used the invariance property (2.19), and likewise for $L(L_{0}(b\overline{u}))=b\overline{u}$. ∎
Proposition 3.25.
Let $u\in W$ be an element such that $u(\alpha)=\alpha$ for some simple root $\alpha$. Then $u$ is not a solution of Problem 1.1.
Proof.
Set $x(t):=x_{\alpha}(t)$ and $y(t):=x_{-\alpha}(t)$, $t\in\mathbb{C}$. Decompose a generic element $b\in\mathcal{B}_{+}$ as $b=ax_{\alpha}(t)n^{\prime}$ where $n^{\prime}\in\mathcal{N}_{+}$ is the product of $x_{\beta}(t_{\beta})$ for $\beta\in\Pi_{+}\setminus\{\alpha\}$. Let $r\in\mathbb{C}$ be a number such that for a given representative $\overline{u}$ of $u$ and any $t\in\mathbb{C}$, $\overline{u}y(t)\overline{u}^{-1}=y(rt)$. Set
$n:=y(c),\ \ c:=\frac{ra^{-\alpha}-1}{rt}.$ (3.56)
We claim that $nb\overline{u}n\overline{u}^{-1}\in\mathcal{B}_{+}$. Indeed, the inclusion equation is given by
$y(-c)ax(t)n^{\prime}y(rc)\in\mathcal{B}_{+};$ (3.57)
commuting $n^{\prime}$ with $y(rc)$ (here we use the fact that $\alpha$ does not appear in $n^{\prime}$) and $y(-c)$ with $a$, the equation is equivalent to
$y(-a^{-\alpha}c)ax(t)y(rc).$ (3.58)
Applying the relation (3.26), we can further rewrite the equation as
$y(-a^{-\alpha}c)y\left(\frac{rc}{1+rtc}\right)\in\mathcal{B}_{+}$ (3.59)
or
$-a^{\alpha}c+\frac{rc}{1+rtc}=0.$ (3.60)
Now it is a matter of direct check that $c$ given by equation (3.56) is a solution of the above equation. By Lemma 3.24, $u$ is not a solution of Problem 1.1. ∎
Proposition 3.26.
Assume there is a cycle in $\Gamma(u)$ that consists of a collection of pairwise orthogonal simple roots. Then $u$ is not a solution of Problem 1.1 and is not rational.
Proof.
Let $\Delta^{\prime}:=\{\alpha_{i_{j}}\}_{j\in I}\subseteq\Delta\cap\nu^{0}(u)$ be a collection of simple roots enumerated by the cyclic group $I:=\mathbb{Z}_{m}$, such that $a_{i_{j},i_{j+1}}=0$. We assume that $m>1$ (the case of $m=1$ is covered by Proposition 3.25), and that there is a cycle
$u^{-1}(\alpha_{i_{j-1}})\leq\alpha_{i_{j}},\ \ j\in I.$ (3.61)
Clearly, the above equation implies that $u(\alpha_{i_{j}})=\alpha_{i_{j-1}}$. For a given representative $\overline{u}$ of $u$, let $\{g_{j}\in\mathbb{C}\}_{j\in I}$ be a collection of numbers such that
$\overline{u}y_{i_{j}}(t)\overline{u}^{-1}=y_{i_{j-1}}(g_{j}t),\ \ t\in\mathbb{C}.$ (3.62)
Decompose a generic element $b\in\mathcal{B}_{+}$ as
$b=a\left[\prod_{j\in I}x_{i_{j}}(t_{j})\right]n^{\prime}$ (3.63)
where $n^{\prime}$ is the product of $x_{\beta}(t_{\beta})$ for $\beta\in\Pi_{+}\setminus\Delta^{\prime}$, in any order. Writing $n\in\mathcal{N}_{-}$ as
$n=\prod_{j\in I}y_{i_{j}}(c_{j}),$ (3.64)
the objective is to use Lemma 3.24 and to show that the inclusion equation $n^{-1}b\overline{u}n\overline{u}^{-1}\in\mathcal{B}_{+}$ has a nontrivial solution in terms of $c_{j}$, $j\in I$. Indeed, we see that
$\begin{split}n^{-1}b\overline{u}n\overline{u}^{-1}=&\left[\prod_{j\in I}y_{i_{j}}(-c_{j})\right]a\left[\prod_{j\in I}x_{i_{j}}(t_{j})\right]n^{\prime}\prod_{j\in I}y_{i_{j-1}}(g_{j}c_{j})=\\ =&a\left[\prod_{j\in I}y_{i_{j}}(-a^{\alpha_{i_{j}}}c_{j})\right]\left[\prod_{j\in I}x_{i_{j}}(t_{j})\right]\left[\prod_{j\in I}y_{i_{j-1}}(g_{j}c_{j})\right]n^{\prime\prime},\end{split}$ (3.65)
where $n^{{\prime\prime}}\in\mathcal{N}_{+}$ is some element. Since the simple roots in $\Delta^{\prime}$ are pairwise orthogonal, the inclusion equation is equivalent to the equation
$\prod_{j\in I}y_{i_{j}}(-a^{\alpha_{i_{j}}}c_{j})x_{i_{j}}(t_{j})y_{i_{j}}(g_{j+1}c_{j+1})\in\mathcal{B}_{+}.$ (3.66)
Commuting $x_{i_{j}}$ with $y_{i_{j}}$ via (3.26), the above equation is equivalent to
$\prod_{j\in I}y_{i_{j}}(-a^{\alpha_{i_{j}}}c_{j})y_{i_{j}}\left(\frac{g_{j+1}c_{j+1}}{1+t_{j}g_{j+1}c_{j+1}}\right)\in\mathcal{B}_{+}.$ (3.67)
Setting $a_{j}:=a^{\alpha_{i_{j}}}$, $j\in I$, the above equation is equivalent to the system of equations
$-a_{j}c_{j}+\frac{g_{j+1}c_{j+1}}{1+t_{j}g_{j+1}c_{j+1}}=0,\ \ j\in I.$ (3.68)
The nontrivial solution to the above system is given as follows. Set
$D_{j}:=t_{j}\prod_{k\neq j}g_{k}+t_{j+1}a_{j+1}\prod_{k\notin\{j,j+1\}}g_{k}+t_{j+2}a_{j+1}a_{j+2}\prod_{k\notin\{j,j+1,j+2\}}g_{k}+\cdots+t_{j-1}\prod_{k\neq j}a_{k}$ (3.69)
where we take advantage of the fact that $I=\mathbb{Z}_{m}$, a cyclic group; then the nontrivial solution for a generic $b$ is given by
$c_{j}=\frac{\prod_{k\in I}g_{k}-\prod_{k\in I}a_{k}}{a_{j}g_{j}D_{j}},\ \ j\in I.$ (3.70)
By Lemma 3.24, we conclude that $u$ is not a solution of Problem 1.1. ∎
Proposition 3.27.
Let $u\in W$ be an element with the following properties: 1) there are simple roots $\alpha_{i}$ and $\alpha_{j}$ with $a_{ij}=a_{ji}=-1$ such that $\alpha_{i},\alpha_{i}+\alpha_{j}\in\nu^{0}(u)$; 2) there is a $2$-cycle on the vertices $\alpha_{i}$ and $\alpha_{i}+\alpha_{j}$ in $\Gamma(u)$. Then $u$ is not a solution of Problem 1.1.
Proof.
Under the above conditions, the existence of the $2$-cycle implies that $u$ acts as a simple reflection $s_{j}$ on the $A_{2}$ root system spanned by $\alpha_{i}$ and $\alpha_{j}$; that is,
$u(\alpha_{i})=u^{-1}(\alpha_{i})=\alpha_{i}+\alpha_{j};$ (3.71)
$u(\alpha_{j})=-\alpha_{j}.$ (3.72)
In other words, $s_{j}u(\alpha_{i})=\alpha_{i}$ and $s_{j}u(\alpha_{j})=\alpha_{j}$. Choose $\overline{s_{i}}:=\dot{s}_{i}$, as in (3.34). Set
$n:=y_{i}(c_{1})y_{j}(c_{2})y_{i}(c_{3}),\ \ c_{1},c_{2},c_{3}\in\mathbb{C},$ (3.73)
and let $r_{i},r_{j}\in\mathbb{C}$ be the numbers such that for any $t\in\mathbb{C}$,
$\overline{s_{j}u}y_{i}(t)\overline{s_{j}u}^{-1}=y_{i}(r_{i}t),$ (3.74)
$\overline{s_{j}u}y_{j}(t)\overline{s_{j}u}^{-1}=y_{j}(r_{j}t).$ (3.75)
Therefore,
$\overline{s_{j}u}n\overline{s_{j}u}^{-1}=y_{i}(r_{i}c_{1})y_{j}(r_{j}c_{2})y_{i}(r_{i}c_{3});$ (3.76)
conjugating both sides by $\overline{s_{j}}^{-1}(\cdot)\overline{s_{j}}$, we see that
$\overline{u}n\overline{u}^{-1}=\overline{s_{j}}^{-1}y_{i}(r_{i}c_{1})y_{j}(r_{j}c_{2})y_{i}(r_{i}c_{3})\overline{s_{j}}.$ (3.77)
The right-hand side can be easily computed in $\operatorname{SL}_{3}(\mathbb{C})$, and in the end, we arrive at
$\overline{u}n\overline{u}^{-1}=y_{j}\left(-\frac{c_{1}+c_{3}}{r_{j}c_{2}c_{3}}\right)y_{i}(r_{i}r_{j}c_{2}c_{3})y_{j}\left(\frac{c_{1}+c_{3}}{r_{j}c_{2}c_{3}}\right)x_{j}(-r_{j}c_{2}).$ (3.78)
Set
$w_{1}:=w_{1}(c_{1},c_{2},c_{3}):=-\frac{c_{1}+c_{3}}{r_{j}c_{2}c_{3}};$ (3.79)
$w_{2}:=w_{2}(c_{1},c_{2},c_{3}):=r_{i}r_{j}c_{2}c_{3}.$ (3.80)
Decompose a generic element $b\in\mathcal{B}_{+}$ as
$b=ax_{i}(t_{1})x_{j}(t_{2})x_{i}(t_{1})n^{\prime}$ (3.81)
where $n^{\prime}$ is the product of $x_{\beta}(t_{\beta})$, $\beta\in\Pi_{+}\setminus\{\alpha_{i},\alpha_{j},\alpha_{i}+\alpha_{j}\}$ taken in any order. Performing algebraic manipulations similar to the ones in the proof of Proposition 3.26, the inclusion equation $n^{-1}b\overline{u}n\overline{u}^{-1}\in\mathcal{B}_{+}$ is equivalent to the equation
$y_{i}(-a^{\alpha_{i}}c_{3})y_{j}(-a^{\alpha_{j}}c_{2})y_{i}(-a^{\alpha_{i}}c_{1})x_{i}(t_{1})x_{j}(t_{2})x_{i}(t_{1})y_{i}(w_{1})y_{j}(w_{2})y_{i}(-w_{1})\in\mathcal{B}_{+}.$ (3.82)
This equation can be solved for $(c_{1},c_{2},c_{3})$ in $\operatorname{SL}_{3}(\mathbb{C})$. Let us state one of the nontrivial solutions; set
$\Delta:=a^{2\alpha_{j}}t_{2}^{2}-4r_{j}a^{\alpha_{j}}.$ (3.83)
Then
$c_{1}=\frac{1}{a^{\alpha_{i}}(t_{1}+t_{3})};$ (3.84)
$c_{2}=\frac{(t_{1}+t_{3})(2r_{i}r_{j}+a^{\alpha_{i}}\Delta^{1/2}+a^{\alpha_{i}}a^{\alpha_{j}}t_{2})}{2a^{\alpha_{j}}r_{j}(a^{\alpha_{i}}(t_{1}+t_{3})+r_{i}t_{2}t_{3})};$ (3.85)
$c_{3}=\frac{-2a^{\alpha_{i}}r_{i}(t_{1}+t_{3}+a^{\alpha_{j}}t_{1}t_{2}^{2}t_{3}-r_{j}(t_{1}+t_{3})^{2})}{a^{\alpha_{i}}(t_{1}+t_{3})+t_{3}\Delta^{1/2}(r_{i}t_{2}t_{1}+1)+a^{\alpha_{j}}t_{2}(t_{1}-t_{3})}.$ (3.86)
Thus the statement holds. ∎
3.6 Proof of Theorem 2.4
In this subsection, we provide a proof of Theorem 2.4. One direction of the theorem is contained in Corollary 3.20: if $u\in W$ is rational, then $u$ is a solution of Problem 1.1. For the second part of the theorem, the case of $\ell(u)=\ell(w_{0})-1$ is treated in Proposition 3.28, and the case of $\ell(u)=\ell(w_{0})-2$ is treated in Proposition 3.29.
Proposition 3.28.
An element $s_{i}w_{0}$ is rational if and only if $s_{i}w_{0}$ is a solution of Problem 1.1, if and only if $w_{0}(\alpha_{i})\neq-\alpha_{i}$;
Proof.
We see that the rationality graph $\Gamma(s_{i}w_{0})$ consists of a single vertex $\alpha_{i}$; hence, $s_{i}w_{0}$ is not rational if and only if there is a loop at $\alpha_{i}$, if and only if $(s_{i}w_{0})^{-1}(\alpha_{i})=\alpha_{i}$, if and only if $w_{0}(\alpha_{i})=-\alpha_{i}$. By Proposition 3.25, if $s_{i}w_{0}(\alpha_{i})=\alpha_{i}$, then $s_{i}w_{0}$ is not a solution; by Corollary 3.20, if $s_{i}w_{0}$ is not a solution, it is not rational, and by the above, $w_{0}(\alpha_{i})=-\alpha_{i}$. ∎
Proposition 3.29.
For $i\neq j$, an element $s_{i}s_{j}w_{0}$ is rational if and only if $s_{i}s_{j}w_{0}$ is a solution of Problem 1.1, if and only if the following conditions are satisfied: 1) $w_{0}(\alpha_{j})\notin\{-\alpha_{i},-\alpha_{j}\}$; 2) If $a_{ij}=0$, then $w_{0}(-\alpha_{i})\neq\alpha_{i}$.
Proof.
We see that the vertex set of $\Gamma(s_{i}s_{j}w_{0})$ is given by
$\nu^{0}(s_{i}s_{j}w_{0})=\{\alpha_{i},s_{i}(\alpha_{j})\}.$ (3.87)
If $s_{i}s_{j}w_{0}$ is not rational, then there is either a loop at $\alpha_{i}$, a loop at $s_{i}(\alpha_{j})$, or a $2$-cycle between the two vertices. A simple computation shows that the loops and the $2$-cycle appear under the following circumstances:
Loop at $\alpha_{i}$.
That is,
$(s_{i}s_{j}w_{0})^{-1}(\alpha_{i})\leq\alpha_{i},$ (3.88)
which is equivalent to the statement $w_{0}(\alpha_{i})=-\alpha_{i}$ and $a_{ij}=0$. Under these conditions, $s_{i}s_{j}w_{0}(\alpha_{i})=\alpha_{i}$; hence, by Proposition 3.25, $s_{i}s_{j}w_{0}$ is not a solution.
Loop at $s_{i}(\alpha_{j})$.
That is, the following equation holds:
$(s_{i}s_{j}w_{0})^{-1}(s_{i}(\alpha_{j}))\leq s_{i}(\alpha_{j}).$ (3.89)
Then either a) $w_{0}(\alpha_{j})=-\alpha_{j}$, or b) $w_{0}(\alpha_{j})=-\alpha_{i}$ and $a_{ij}\neq 0$. Indeed, using equation (3.20), we can expand (3.89) as
$w_{0}(-\alpha_{j})\leq\alpha_{j}-a_{ij}\alpha_{i},$ (3.90)
and then it’s easy to see the equivalence. In case of a), $s_{j}w_{0}$ is not a solution by Proposition 3.28, and since $s_{j}w_{0}(\alpha_{i})\in\Pi_{-}$, by Corollary 3.22, $s_{i}s_{j}w_{0}$ is not a solution. If b) is the case, then $\Gamma(s_{i}s_{j}w_{0})$ also has a $2$-cycle (see the next item).
A $2$-cycle.
This is the case if and only if $w_{0}(\alpha_{j})=-\alpha_{i}$. Indeed, there is a $2$-cycle if and only if the following system of inequalities is satisfied:
$(s_{i}s_{j}w_{0})^{-1}(\alpha_{i})\leq s_{i}(\alpha_{j});$ (3.91)
$(s_{i}s_{j}w_{0})^{-1}(s_{i}(\alpha_{j}))\leq\alpha_{i}.$ (3.92)
We see that the second equation is satisfied if and only if $w_{0}(\alpha_{j})=-\alpha_{i}$, and in this case, the Lie type of $W$ is simply laced, so $a_{ij}=-a_{ji}$; hence, $-w_{0}(s_{j}(\alpha_{i}))=s_{i}(\alpha_{j})$, and the first equation is satisfied. Now, if $a_{ij}=0$, then there is a cycle between two orthogonal simple roots, and by Proposition 3.26, $s_{i}s_{j}w_{0}$ is not a solution. If $a_{ij}=-1$, then $s_{i}s_{j}w_{0}$ is not a solution by Proposition 3.27.
Thus the statement holds. ∎
Corollary 3.30.
Assume that the Lie type is such that $w_{0}=-\operatorname{id}$, and let $u\in W$ be any element of length $\ell(u)\geq\ell(w_{0})-2$. Then $u$ is a solution of Problem 1.1 if and only if $u$ is rational, if and only if $u=w_{0}$.
Proof.
Indeed, these are the Lie types precisely in which $w_{0}(\alpha_{i})=-\alpha_{i}$ for every simple root $\alpha_{i}$. The corollary follows from Proposition 3.29. ∎
Corollary 3.31.
In type $A_{2}$, an element $u\in W$ is a solution of Problem 1.1 if and only if $u$ is rational. In particular, the set of solutions of Problem 1.1 in type $A_{2}$ is given by
$\{w_{0}=s_{1}s_{2}s_{1},C:=s_{1}s_{2},C^{-1}=s_{2}s_{1}\}.$ (3.93)
4 Combinatorial results
In this section, we provide proofs of the results from Section 2.2. Throughout the section, we assume the root system of $G$ is indecomposable.
4.1 Proof of Theorem 2.11
In this subsection, we prove Theorem 2.11. We split the statement of the theorem into three separate parts: Proposition 4.5, Corollary 4.6 and Proposition 4.7.
Lemma 4.1.
For any $u\in W$ and any simple root $\alpha$,
$\nu^{0}(s_{\alpha}u)=s_{\alpha}(\nu^{0}(u)\setminus\{\alpha\})\cup(s_{\alpha}u(\Pi_{+})\cap\{\alpha\}).$ (4.1)
Proof.
Indeed,
$\begin{split}\nu^{0}(s_{\alpha}u)=s_{\alpha}u(\Pi_{+})\cap\Pi_{+}=&s_{\alpha}\left(u(\Pi_{+})\cap(\Pi_{+}\setminus\{\alpha\}\cup\{-\alpha\})\right)=\\ =&s_{\alpha}(\nu^{0}(u)\setminus\{\alpha\})\cup(s_{\alpha}u(\Pi_{+})\cap\{\alpha\}).\end{split}$ (4.2)
Thus the equation holds. ∎
For the next lemma, given a simple root $\alpha\in\Delta$ and a root $\beta\in\Pi$, denote by $c_{\alpha}(\beta)$ the coefficient of $\beta$ at $\alpha$ in its expansion of simple roots; that is, $\beta=\sum_{\alpha\in\Delta}c_{\alpha}(\beta)\alpha$.
Lemma 4.2.
Let $u\in W$ be a rational Weyl group element such that $u\neq w_{0}$. Then there exists a simple root $\alpha$ such that $u^{-1}(\alpha)>0$ and $u(\alpha)<0$.
Proof.
Since $u\neq w_{0}$ and $u$ is rational, there is $m\geq 1$ such that $\nu^{m}(u)\neq\emptyset$ and $\nu^{m+1}(u)=\emptyset$. Pick any $\beta\in\nu^{m}(u)$; since $u^{-1}(\beta)>0$, there is a simple root $\alpha$ such that $\alpha\leq\beta$ and $u^{-1}(\alpha)>0$. Hence $\alpha\in\operatorname{Adj}\nu^{m}(u)$, and since $\nu^{m+1}(u)=\emptyset$, we see that $u(\alpha)<0$. Thus $\alpha$ is the required root. ∎
Lemma 4.3.
Let $\gamma\in\Pi_{+}$, $\alpha\in\Delta$ and $c>0$ be such that $\gamma-c\alpha\in\Pi_{+}$. Let $\beta\in\Pi_{+}$ be such that $\beta\leq\gamma$, $c_{\alpha}(\beta)=c_{\alpha}(\gamma)$ and $\beta\neq\alpha$. Then $\beta-c\alpha\in\Pi_{+}$.
Proof.
From the given conditions, observe that $\langle\beta,\alpha^{\vee}\rangle\geq\langle\gamma,\alpha^{\vee}\rangle$. In the simply-laced case, we see that $c=1$, and since $\beta\neq\alpha$, the above equation implies $\langle\beta,\alpha^{\vee}\rangle=1$; hence, $s_{\alpha}(\beta)=\beta-c\alpha\in\Pi_{+}$.
For the non-simply-laced case, we consider each type separately. For $G_{2}$, there are no roots $\gamma$ and $\beta$ that satisfy the conditions of the lemma; for $F_{4}$, one can observe the root poset directly. Types $B_{r}$ and $C_{r}$ are similar in treatment, so let us only verify the type $B_{r}$.
In type $B_{r}$, enumerate the simple roots so that $\alpha_{r}$ is the short one. Since the statement holds for simply-laced types, we assume that $c_{\alpha_{r}}(\gamma)\neq 0$. If $\gamma=\sum_{k=i}^{j-1}\alpha_{k}+2\sum_{k=j}^{r}\alpha_{k}$ for $1\leq i<j\leq r$, then $\alpha\in\{\alpha_{i},\alpha_{j}\}$. If $\alpha=\alpha_{i}$, then $\beta=\sum_{k=i}^{t-1}\alpha_{k}+2\delta\sum_{k=t}^{r}\alpha_{k}$ for some $t\in[j,r+1]$ and $\delta\in\{0,1\}$; hence, $\beta-\alpha_{i}\in\Pi_{+}$. If $\alpha=\alpha_{j}$, then $\beta=\sum_{k=p}^{j-1}\alpha_{k}+2\sum_{k=j}^{r}\alpha_{k}$ for some $p\geq i$, so $\beta-c\alpha_{i}\in\Pi_{+}$ (note: $c\in\{1,2\}$ for $j=r$; if $j<r$, then $c=1$). Thus the statement holds. ∎
Lemma 4.4.
Let $u\in W$ be a rational Weyl group element such that $u\neq w_{0}$. For any simple root $\alpha$ such that $u^{-1}(\alpha)>0$ and $u(\alpha)<0$, the element $s_{\alpha}u$ is rational.
Proof.
Assume on the contrary that the element $s_{\alpha}u$ is not rational. Then the rationality graph $\Gamma(u)$ contains a cycle $\{\theta_{i}\in\nu^{0}(u)\}_{i\in I}$ where we index the elements by $I:=\mathbb{Z}_{m}$, so that
$(s_{\alpha}u)^{-1}(\theta_{i-1})\leq\theta_{i},\ i\in I.$ (4.3)
Since $u^{-1}(\alpha)>0$, from equation (4.1) we see that
$\nu^{0}(s_{\alpha}u)=s_{\alpha}(\nu^{0}(u)\setminus\{\alpha\});$ (4.4)
therefore, we can write $\theta_{i}=s_{\alpha}\gamma_{i}$ for $\gamma_{i}\in\nu^{0}(u)\setminus\{\alpha\}$, $i\in I$, and equation (4.3) becomes
$u^{-1}(\gamma_{i-1})\leq s_{\alpha}\gamma_{i},\ \ i\in I.$ (4.5)
Define the following subsets of indices:
$I_{-}:=\{i\in I\ |\ u^{-1}(\gamma_{i-1})\leq\gamma_{i}\};$ (4.6)
$I_{+}:=I\setminus I_{-}.$ (4.7)
Observe that if $s_{\alpha}\gamma_{i}=\gamma_{i}-c_{i}\alpha$ for some $c_{i}\geq 0$, then $i\in I_{-}$; hence, if $i\in I_{+}$, then necessarily $s_{\alpha}\gamma_{i}=\gamma_{i}+c_{i}\alpha$ for some $c_{i}>0$. Therefore, there is a number $c_{i}^{\prime}$ such that $0<c_{i}^{\prime}\leq c_{i}$ and $\lambda_{i}\in\mathbb{Z}(\Delta\setminus\{\alpha\})$ such that
$u^{-1}(\gamma_{i-1})+\lambda_{i}=\gamma_{i}+c_{i}^{\prime}\alpha.$ (4.8)
By Lemma 4.3, since $\gamma_{i}-\lambda_{i}+c_{i}^{\prime}\alpha\in\Pi_{+}$ and $\gamma_{i}+c_{i}^{\prime}\alpha\in\Pi_{+}$, we see that [^4] To align with the statement of Lemma 4.3, $\beta:=\gamma_{i}-\lambda_{i}+c_{i}^{\prime}\alpha$ and $\gamma:=\gamma_{i}+c_{i}^{\prime}\alpha$; so, $\gamma-c_{i}^{\prime}\alpha\in\Pi_{+}$. The case $\beta=\alpha$ is excluded due to the assumption $u(\alpha)<0$ (if $\beta=\alpha$, then $\gamma_{i-1}=u(\alpha)<0$, which is a contradiction). $\gamma_{i}-\lambda_{i}\in\Pi_{+}$. Applying $u$ to the above equation, we see that
$\gamma_{i-1}=u(\gamma_{i}-\lambda_{i})+c_{i}^{\prime}u(\alpha),$ (4.9)
and since $u(\alpha)<0$, we obtain the equation
$\gamma_{i-1}<u(\gamma_{i}-\lambda_{i}).$ (4.10)
In particular, $u(\gamma_{i}-\lambda_{i})\in\nu^{0}(u)$. We now exhibit paths in $\Gamma(u)$ which, when combined together, yield a cycle in $\Gamma(u)$.
Case $i,i+1\in I_{-}$.
There is an arrow
$\gamma_{i}\rightarrow\gamma_{i+1}.$ (4.11)
Case $i\in I_{-}$, $i+1\in I_{+}$.
It follows from equation (4.10) that
$u^{-1}(\gamma_{i-1})\leq\gamma_{i}<u(\gamma_{i+1}-\lambda_{i+1}),$ (4.12)
and therefore, there is a path
$\gamma_{i-1}\rightarrow u(\gamma_{i+1}-\lambda_{i+1})\rightarrow\gamma_{i+1}.$ (4.13)
Case $i,i+1\in I_{+}$.
Due to equation (4.10), there are inequalities
$u^{-1}(u(\gamma_{i}-\lambda_{i}))=\gamma_{i}-\lambda_{i}\leq\gamma_{i}<u(\gamma_{i+1}-\lambda_{i+1}),$ (4.14)
and therefore, there is a path
$u(\gamma_{i}-\lambda_{i})\rightarrow u(\gamma_{i+1}-\lambda_{i+1})\rightarrow\gamma_{i+1}.$ (4.15)
Case $i\in I_{+}$, $i+1\in I_{-}$.
There is an arrow
$u(\gamma_{i}-\lambda_{i})\rightarrow\gamma_{i+1}.$ (4.16)
We thus conclude that $s_{\alpha}u$ is rational. ∎
Proposition 4.5.
The rationality graph $\Gamma(W)$ is connected.
Proof.
Since $w_{0}$ is a rational Weyl group element, it suffices to prove there is a path in $\Gamma(W)$ between any rational $u\in W$ and $w_{0}$. Write $u=vw_{0}$ for some $v\in W$, and proceed by induction on $\ell(v)$. If $\ell(v)=0$, then $u=w_{0}$ is rational. For a fixed $k>0$, assume there is a path in $\Gamma(W)$ between $u$ with $\ell(v)=k-1$ and $w_{0}$. By Lemma 4.2 and Lemma 4.4, there exists a simple root $\alpha$ such that $v^{-1}(\alpha)<0$ and $s_{\alpha}u$ is a rational element. But then, $\ell(s_{\alpha}v)=k-1$, and by the assumption, there is a path between $s_{\alpha}u$ and $w_{0}$; thus, there is a path between $u$ and $w_{0}$. ∎
Corollary 4.6.
The rationality graph $\Gamma(W)$ has more than one vertex only in the following Lie types: $A_{r}$ for $r\geq 2$, $D_{r}$ for $r$ odd, and $E_{6}$.
Proof.
Indeed, by Proposition 4.5, the rationality graph is connected. However, for $\alpha\in\Delta$, $s_{\alpha}w_{0}$ is rational if and only if $w_{0}(\alpha)\neq-\alpha$. Indeed, if $w_{0}(\alpha)=-\alpha$, then $s_{\alpha}w_{0}(\alpha)=\alpha$, and by Proposition 3.5, $s_{\alpha}w_{0}$ is not rational. This implies that $w_{0}$ is connected to a rational Weyl group element in $\Gamma(W)$ if and only if $w_{0}$ induces a nontrivial automorphism of the corresponding Dynkin diagram. It is so precisely in the Lie types from the statement of the corollary. ∎
For the next proposition, let $\epsilon:\Delta\rightarrow\Delta$ be an automorphism of the Dynkin diagram such that $w_{0}(\alpha)=-\epsilon(\alpha)$. Then $\epsilon$ induces an involutive automorphism of the corresponding Weyl group $W$ via $\epsilon(s_{\alpha}):=s_{\epsilon(\alpha)}$.
Proposition 4.7.
For any element $u\in W$, $u$ is rational if and only if $\epsilon(u)$ is rational.
Proof.
Equivalently, for $u\in W$, set $u=vw_{0}$. Since $\epsilon(w_{0})=w_{0}$, we see that $u$ is rational if and only if $\epsilon(v)w_{0}$ is rational. Now the result follows from the observation that the elements of $\nu^{0}(u)$ can be determined via a reduced decomposition of $v$ (see Remark 3.8). ∎
4.2 Proof of Proposition 2.12 (Coxeter elements)
In this subsection, we prove Proposition 2.12, which states that the only rational Coxeter elements are in type $A_{r}$, and they are of the form $C:=s_{1}\cdots s_{r}$ and $C^{-1}$. We split the proposition into three separate statement: for type $A_{r}$, Proposition 4.9 and Proposition 4.10; for type $D_{r}$, Proposition 4.12; for type $E_{6}$, the statement can be verified with computer software (see Proposition 4.13). For all the other types, the statement is a consequence of Corollary 4.6.
For the next results, we model the root system of type $A_{r}$ in an $(r+1)$-dimensional Euclidean space $E^{r+1}$ endowed with an orthonormal basis $e_{1},\ldots,e_{r+1}$. The set of positive roots is given by
[…]
and the set of simple roots is given by
[…]
Each simple reflection $s_{i}$ acts on $E^{r+1}$ by permuting the vectors $e_{i}$ and $e_{i+1}$ and fixing $e_{j}$ for $j\notin\{i,i+1\}$. The Coxeter element $C$ acts by
$C(e_{i})=\begin{cases}e_{i+1}&i\in\{1,\ldots,r\}\\ e_{1}&i={r+1}.\end{cases}$ (4.19)
For the next lemma, set
$C_{1}:=s_{1}s_{2}\cdots s_{r}=C;$
$C_{2}:=s_{2}s_{3}\cdots s_{r}s_{1};$
$\vdots$
$C_{m}:=s_{m}s_{m+1}\cdots s_{r}s_{m-1}s_{m-2}\cdots s_{1};$
$\vdots$
$C_{r-1}:=s_{r-1}s_{r}s_{r-2}s_{r-3}\cdots s_{1}$
$C_{r}:=s_{r}s_{r-1}\cdots s_{1}=C^{-1}.$
Lemma 4.8.
Let $c$ be a Coxeter element in type $A_{r}$ and $\theta$ be the highest root. Then $c(\theta)>0$ if and only if $c\notin\{C_{1},C_{2},\ldots,C_{r}\}$. Moreover, $C_{m}(\theta)=-\alpha_{m}$, $m\in\{1,\ldots,r\}$.
Proof.
Note that the highest root $\theta$ is given by $\theta=e_{1}-e_{r+1}$, and a direct check shows that $C_{m}(\theta)=-\alpha_{m}$. Let us embed the Euclidean space $E^{r}$ into $E^{r+1}$ as the orthogonal complement of the vector $e_{r+1}$. We proceed by induction on $r$, with cases $r=1$ and $r=2$ being evident. A Coxeter element $c$ can be written as $s_{i_{1}}\cdots s_{i_{r-1}}s_{r}$ or as $c=s_{r}s_{i_{1}}\cdots s_{i_{r-1}}$. In the first case, we see that
$s_{i_{1}}\cdots s_{i_{r-1}}s_{r}(e_{1}-e_{r+1})=s_{i_{1}}\cdots s_{i_{r-1}}(e_{1}-e_{r}),$ (4.20)
and hence the statement follows from the case of $A_{r-1}$. In the second case, we see that $c(\theta)=s_{r}(e_{i}-e_{r+1})$ for some $i\in\{1,\ldots,r\}$. If $i<r$, then $c(\theta)>0$; if $i=r$, then $c(\theta)=-\alpha_{r}$, and in this case, $c=s_{r}s_{r-1}\cdots s_{1}$. ∎
Proposition 4.9.
In type $A_{r}$, a Coxeter element $c\in W$ is rational if and only if $c\in\{C,C^{-1}\}$.
Proof.
By Lemma 4.8, we see that if $c\notin\{C_{1}^{-1},C_{2}^{-1},\ldots,C_{r}^{-1}\}$, then $c^{-1}(\theta)>0$, and therefore, $\theta\in\nu^{0}(c)$; by Proposition 3.4, $c$ is not rational.
Assume that $c=C_{m}^{-1}$ for some $m\in\{2,\ldots,r-1\}$. Observe that there is a $2$-cycle in $\Gamma(c)$:
$(c^{-1})(e_{1}-e_{m+1})=e_{m+1}-e_{m+2}\leq e_{m}-e_{r+1};$ (4.21)
$(c^{-1})(e_{m}-e_{r+1})=e_{m-1}-e_{m}\leq e_{1}-e_{m+1};$ (4.22)
hence, $c$ is not rational.
It remains to show that if $c\in\{C,C^{-1}\}$, then $c$ is rational. Observe that
$\nu^{k-1}(C)=\{e_{i}-e_{j}\ |\ k<i<j\leq r+1\},\ \ k\geq 1.$ (4.23)
Evidently, $\nu(C)=\emptyset$. To show that $C^{-1}$ is rational, observe that $\epsilon(C)=C^{-1}$ (the $\mathbb{Z}_{2}$-symmetry of $\Gamma(W)$, see Proposition 4.7), hence $C^{-1}$ is rational. ∎
Proposition 4.10.
In type $A_{r}$, the Coxeter elements $C$ and $C^{-1}$ are rational and have valency $1$ in the graph $\Gamma(W)$. More precisely, $C$ is connected only to the rational element $s_{r}C$, and $C^{-1}$ is connected only to the rational element $s_{1}C^{-1}$.
Proof.
The Coxeter elements $C$ and $C^{-1}$ are rational by Proposition 4.9. We see that $C(\alpha_{r})<0$ and $C^{-1}(\alpha_{r})>0$; hence, by Lemma 4.4, the element $s_{r}C$ is rational. Likewise, by the same lemma, the element $s_{1}C^{-1}$ is rational.
Let us show that the valency of $C$ in $\Gamma(W)$ is equal to $1$. Consider the element $s_{i}C$ for $i\in\{1,\ldots,r-1\}$. We see that
$s_{i}C(\alpha_{i})=s_{i}(\alpha_{i+1})=\alpha_{i}+\alpha_{i+1}\geq\alpha_{i};$ (4.24)
therefore, by Proposition 3.5, $s_{i}C$ is not rational, and thus the valency of $C$ is $1$. By the $\mathbb{Z}_{2}$-symmetry of $\Gamma(W)$ (see Proposition 4.7), the valency of $C^{-1}$ is also $1$. ∎
For the next results, we model the root system of type $D_{r}$ in an $r$-dimensional Euclidean space $E^{r}$ endowed with an orthonormal basis $e_{1},\ldots,e_{r}$. Then the set of positive roots is given by
[…]
and the set of simple roots is given by
[…]
In this model, each simple reflection $s_{i}$ for $i\in\{1,\ldots,r-1\}$ interchanges $e_{i}$ and $e_{i+1}$, and leaves the other basis vectors fixed; for $s_{r}$, $s_{r}(e_{r-1})=-e_{r}$.
Lemma 4.11.
Let $\theta$ be the highest root vector in type $D_{r}$. For any Coxeter element $c$, $c(\theta)>0$.
Proof.
Note that the highest root $\theta$ in type $D_{r}$, in the given Euclidean model, is $\theta=e_{1}+e_{2}$. Consider two cases based on the position of the simple reflection $s_{r}$ in a reduced decomposition of $c$.
Case 1.
Assume that $c=s_{i_{1}}s_{i_{2}}\cdots s_{i_{r-1}}s_{r}$. Since $s_{r}(\theta)=\theta$, and since no other simple reflection can change the sign of $\theta$, we see that $c(\theta)>0$.
Case 2.
Assume that $c=s_{r}s_{i_{1}}\cdots s_{i_{r-1}}$. Then $s_{i_{1}}\cdots s_{i_{r-1}}(\theta)=e_{i}+e_{j}$ for some $i\neq j$. Then $c(\theta)<0$ if and only if $i=r-1$ and $j=r$. However, $s_{i_{1}}\cdots s_{i_{r-1}}(e_{1})=e_{r-1}$ if and only if $(i_{1},\ldots,i_{r-2},i_{r-1})=(r-1,r-2,\ldots,1)$, but in this case, $s_{r-1}s_{r-2}\cdots s_{1}(e_{2})=e_{2}$.
Thus $c(\theta)>0$ for any Coxeter element $c$ in type $D_{r}$. ∎
Proposition 4.12.
In type $D_{r}$, no Coxeter element is rational.
Proof.
The statement is a direct consequence of Lemma 4.11 and Proposition 3.4. ∎
Proposition 4.13.
In type $E_{6}$, no Coxeter element is rational.
Proof.
This can be verified with computer software via seeing, as before, that for every Coxeter element $c$ in type $E_{6}$ and the highest root $\theta$, $\theta\in\nu^{0}(c)$. ∎
4.3 Proof of Proposition 2.13 (valency $1$ vertices in $\Gamma(D_{r})$)
In this subsection, we construct a pair of rational elements in type $D_{r}$ of valency $1$ in the rationality graph $\Gamma(D_{r})$. We establish some of its basic properties in Proposition 4.14, and we prove Proposition 2.13 (whose more precise statement is Proposition 4.15).
We represent the root system $D_{r}$ in a Euclidean space $E^{r}$ in the same way as in Section 4.3. Define the element $C$ by its action upon the orthonormal basis $e_{1},\ldots,e_{r}$, as follows:
$C(e_{i})=\begin{cases}-e_{r}&\text{if}\ i=1\\ -e_{i}&\text{if}\ i\in\{2,3,\ldots,r-1\}\\ e_{1}&\text{if}\ i=r.\end{cases}$ (4.27)
Proposition 4.14.
The element $C$ belongs to $W$; moreover, its length is equal to
$\ell(C)=\frac{1}{2}(r(r-1)+(r-2)(r-3)),$ (4.28)
its order is equal to $4$, and its reduced decomposition can be constructed as [^5] The square parentheses carry no extra meaning and are to be interpreted as round parentheses.
$\begin{split}C=&s_{r}(s_{r-2}s_{r-1})[s_{r}](s_{r-3}s_{r-2}s_{r-1})[s_{r}s_{r-2}](s_{r-4}s_{r-3}s_{r-2}s_{r-1})\cdot\\ \cdot&[s_{r}s_{r-2}s_{r-3}](s_{r-5}s_{r-3}s_{r-2}s_{r-1})\cdots[s_{r}s_{r-2}s_{r-3}\cdots s_{3}](s_{1}s_{2}\cdots s_{r-2}s_{r-1}).\end{split}$ (4.29)
Lastly, $\epsilon(C)=C^{-1}$ where $\epsilon$ is the $\mathbb{Z}_{2}$-symmetry of $\Gamma(D_{r})$.
Proof.
The formulas can be verified via a direct computation in the orthonormal basis $e_{1},\ldots,e_{r}$. To verify that $C\cdot\epsilon(C)=\operatorname{id}$, one computes directly the product using the decomposition (4.29). ∎
Proposition 4.15.
In type $D_{r}$ for $r\geq 5$ odd, the elements $C$ and $C^{-1}$ are rational and have valency $1$ in the rationality graph $\Gamma(D_{r})$. More precisely, $C$ is connected only to the rational element $s_{r-1}C$, and $C^{-1}$ is connected only to the rational element $s_{r}C^{-1}$.
Proof.
From the $\nu$-sequence for $C$, we see that
$\nu^{0}(C)=\{e_{1}-e_{i}\ |\ i\in\{2,\ldots,r-1\}\}\cup\{e_{i}-e_{r}\ |\ i\in\{1,\ldots,r-1\}\};$ (4.30)
$\nu^{1}(C)=\{e_{i}-e_{r}\ |\ i\in\{2,\ldots,r-1\}\};$ (4.31)
$\nu^{2}(C)=\emptyset;$ (4.32)
therefore, $C$ is rational. By the $\mathbb{Z}_{2}$-symmetry of $\Gamma(D_{r})$ (see Proposition 4.7), $C^{-1}$ is also rational. We see that
$C(e_{r-1}-e_{r})=-(e_{1}+e_{r-1})<0,\ \ C^{-1}(e_{r-1}-e_{r})=e_{1}-e_{r-1}>0.$ (4.33)
By Lemma 4.4, $s_{r-1}C$ is rational. Likewise, one verifies that $s_{r}C^{-1}$ is rational by the same lemma.
Now, let us show that $s_{i}C$ for $i\in\{1,\ldots,r-2\}\cup\{r\}$ is not rational. For $i\in\{1,\ldots,r-2\}$, we see that
$s_{i}C(e_{i}-e_{i+1})=e_{i}-e_{i+1};$ (4.34)
and for $i=r$, we see that
$s_{r}C(e_{r-1}+e_{r})=s_{r}(e_{1}-e_{r-1})=e_{1}+e_{r}\geq e_{r-1}+e_{r}.$ (4.35)
By Proposition 3.5, $s_{i}C$ for $i\in\{1,\ldots,r-2\}\cup\{r\}$ is not rational. ∎
Appendix A Rationality graphs $\Gamma(A_{3})$, $\Gamma(A_{4})$ and $\Gamma(D_{5})$
(a) Rationality graph $\Gamma(A_{3})$.
(b) Rationality graph $\Gamma(A_{4})$.
(c) Rationality graph $\Gamma(D_{5})$.
Figure 1: Rationality graphs in types $A_{3}$, $A_{4}$ and $D_{5}$. Each label $i$ of an edge is to be interpreted as a simple reflection $s_{i}$.